### v1

# Quasi-Monte Carlo with one categorical variable

Abstract
We study randomized quasi-Monte Carlo (RQMC) estimation of a multivariate integral where one of the variables takes only a finite number of values. This problem arises when the variable of integration is drawn from a mixture distribution as is common in importance sampling and also arises in some recent work on transport maps. We find that when integration error decreases at an RQMC rate that it is then beneficial to oversample the smallest mixture components instead of using a proportional allocation. We also find that for the most accurate RQMC sampling methods, it is advantageous to arrange that our $n=2^{m}$ randomized Sobol’ points split into subsample sizes that are also powers of $2$.
1 Introduction
We study randomized quasi-Monte Carlo (RQMC) integration in a setting where one of the input variables is categorical, taking only some finite number $L\geqslant 2$ of different values. The main motivation for this problem is mixture sampling where the variables of interest come from $L$ different distributions representing $L$ different scenarios. That is $\bm{x}\sim P_{\ell}$ with probability $\alpha_{\ell}>0$ where $\sum_{\ell=1}^{L}\alpha_{\ell}=1$. The distributions $P_{\ell}$ are commonly called mixture components. We call them ‘strata’ to avoid confusion with vector components.
A straightforward approach to this problem is to simply replace a standard Monte Carlo (MC) workflow with RQMC sampling. In such a workflow, one random variable selects stratum $\ell$ with probability $\alpha_{\ell}$. Then subsequent random variables are used to generate $\bm{x}\sim P_{\ell}$. When MC is replaced by RQMC, we find some advantages to sampling the $L$ strata more equally than their nominal proportions $\alpha_{\ell}$ would indicate.
An alternative to the standard MC workflow is to compute $L$ different independently sampled RQMC estimates, one per stratum, and combine them according to their mixture probabilities $\alpha_{\ell}$. That approach misses a potential variance reduction from RQMC sampling when the within-stratum estimates are negatively correlated. Such negative correlation need not happen; we discuss them briefly at the end of the paper.
For some problems the integrand is a simple function within each stratum. A more interesting expression arises in mixture importance sampling. There we estimate $\mu=\int g(\bm{x})p(\bm{x})\mathrm{\,d}\bm{x}$ for a probability density function $p$ by
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\frac{g(\bm{x}_{i})p(\bm{x}_{i})}{\sum_{\ell=1}^{L}\alpha_{\ell}p_{\ell}(\bm{x}_{i})}$ (1)
for $\bm{x}_{i}\sim\sum_{\ell=1}^{L}\alpha_{\ell}p_{\ell}$. By Jensen’s inequality, this has lower Monte Carlo variance than we would get from
$\sum_{\ell=1}^{L}\frac{\alpha_{\ell}}{n_{\ell}}\sum_{i=1}^{n_{\ell}}\frac{g(\bm{x}_{i,\ell})p(\bm{x}_{i,\ell})}{p_{\ell}(\bm{x}_{i,\ell})}.$
working separately within each of the $L$ strata.
RQMC does best under favorable regularity assumptions, such as continuity, on the integrand. The presence of the categorical variable makes the resulting integrand discontinuous. That discontinuity is however axis parallel and so it is a QMC-friendly discontinuity in the terminology of Wang and Tan (2013). In some very regular cases, RQMC can attain a mean squared error (MSE) of $O(n^{-3+\epsilon})$. Such a rate can only be attained over a sequence of special sample sizes that grow at least exponentially (Sobol’, 1998; Owen, 2016). This rate can be attained at powers of $2$ by scrambled Sobol’ points. Even an integrand with a QMC-friendly axis-parallel discrepancy is not generally integrated with an error of this rate. By using an importance sampling (Owen, 2013, Chapter 9) adjustment, we can arrange to obtain $n_{\ell}=2^{m_{\ell}}$ points from $P_{\ell}$ while also having $n=\sum_{\ell=1}^{L}n_{\ell}$ be a power of $2$. This can then give the $O(n_{\ell}^{-3+\epsilon})$ rate within stratum $\ell$ and hence overall.
This paper is organized as follows. Section 2 introduces our notation, gives some background on RQMC, and proves some results about within-stratum discrepancies. Section 3 describes how to allocate samples to strata. In RQMC sampling, it is advantageous to use more nearly equal allocations than those given by the stratum sizes. Section 4 shows how to partition a sample size $n=2^{m}$ into $L$ within-stratum sample sizes $n_{\ell}=2^{m_{\ell}}$. Section 5 establishes some minimax variance results from using equal or nearly equal sample sizes for the strata. Section 6 shows a simple toy example with $L=8$ strata where RQMC with $n_{\ell}$ taken to be powers of two does best. Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg and De Rocquigny (2010) where non-nominal sample sizes do best. Section 8 has some discussion and conclusions. The proof of one of our results is in Appendix A.
We finish this introduction by describing some related work in the literature. Our discussion uses the notation of this paper to describe the findings of the other papers. The key differences between our work and the prior work we found is in how the problem has been formulated, and how the values are sampled.
We use the first component of our RQMC points to select the stratum to sample from. Keller and Grünschloß (2012) used that approach in parallel computation over $L$ processors. Their first component was used to determine which processor $\ell\in\{1,2,\dots,L\}$ got the $i$’th job. The remaining components of their QMC points sampled that job on processor $\ell$. In that setting, $\alpha_{\ell}=1/L$ and it makes sense to choose $L$ to be a power of $2$ at least for Sobol’ sequences.
Cui et al. (2025) consider a target distribution on a bounded hyper rectangle $[\bm{a},\bm{b}]\subset^{d}$. The target probability density function $\pi$ is bounded. They approximate it by a sum of products of hat functions. That is equivalent to a mixture of distributions whose components have independent triangle distributions. They consider adaptive methods to select those triangle densities given an unnormalized $\bar{\pi}\propto\pi$. For a possibly large number of mixture components, they sample each component with $n_{j}\doteq n\alpha_{j}$ input values. They use the first $n_{j}$ points of the same QMC point set for each of the strata. Their integrands $f$ belong to an $L_{p}$ weighted space with a slightly different norm from the usual ones. Their equation (3) defining the norm also uses some sup-norms of absolute mixed partial derivatives. Then sampling on a class of digital nets, they are able to bound the integration error by an expression using the integrand’s norm. They have another bound when the density satisfies a Hölder condition.
We use different RQMC points in each stratum. Our inputs are randomized inputs, and so we expect better error cancellation than we would get using the same inputs for every stratum. Since our analysis is based on variances, we do not require a bounded domain or integrand. We also find an advantage, theoretically and empirically, to diverging from proportional allocation.
Klebanov and Sullivan (2023) describe using transport maps from a $\mathcal{N}(0,I)$ reference distribution to sample a mixture like $P=\sum_{\ell=1}^{L}\alpha_{j}\mathcal{N}(\mu_{j},\Sigma_{j})$. They construct a continuous mapping over $t\in[0,1]$ with $\mathcal{N}(0,I)$ at $t=0$ reaching $P$ at $t=1$ and solve an ordinary differential equation along that path. For $\ell<L$ they take the first $n_{\ell}=\lfloor n\alpha_{\ell}\rfloor$ points of a quasi-Monte Carlo sequence and use those to sample from component $\ell$. Then $n_{L}=n-\sum_{\ell=1}^{L-1}n_{\ell}$. This is different from our strategy of using the first RQMC component to select the stratum which we prefer because it avoids giving two within-stratum samplers the same points. They do not require $n\gg L$. Some of our expressions only require $n>2L$, but as we use asymptotic variances within strata, our results are most relevant for large $n/L$. They consider more general location-scale families than the Gaussian distribution and they can even include some $\alpha_{\ell}<0$ so long as $P$ never has a negative probability density function.
2 Background
Here we introduce notation to describe the different variables used in the integration workflow. Then we define digital nets and discuss their randomizations. For a broader survey of RQMC see L’Ecuyer and Lemieux (2002).
We use $\mathbb{N}$ for the set of positive integers. We use $\bm{1}\{A\}$ for the indicator function that equals $1$ when $A$ holds and is 00 otherwise.
2.1 Integration variables
Our goal is to estimate
$\mu=\sum_{\ell=1}^{L}\alpha_{\ell}\int_{{}^{D}}g(\bm{x})p_{\ell}(\bm{x})\mathrm{\,d}\bm{x}$
where $p_{\ell}$ is the probability density function of $P_{\ell}$ and $g(\cdot)$ is the integrand of interest. We assume that we have functions $\phi_{\ell}$ with $\phi_{\ell}(\bm{u})\sim P_{\ell}$ for $\bm{u}\sim\mathbf{U}(0,1)^{s}$. Many such functions can be obtained using methods like those in Devroye (1986). The dimension $s$ need not equal $D$.
We will use RQMC points $\bm{z}\in[0,1]^{s+1}$. We partition $\bm{z}$ into its first component $v$ that we use to select a stratum $\ell$ and the subsequent $s$ components $\bm{u}$ that we use to sample $\bm{x}$ from $P_{\ell}$. That is, $\bm{z}=(v,u_{1},u_{2},\dots,u_{s})$. For a function $\bm{\ell}:[0,1]\to\{1,2,\dots,L\}$, we will sample the $i$’th point using sample stratum $\bm{\ell}(v_{i})$. The pre-image of each $\ell\in\{1,2,\dots,L\}$ under $\bm{\ell}$ will always be a sub-interval of $[0,1]$. We let $\beta_{\ell}$ be the length of that subinterval. We abbreviate $\bm{\ell}(v_{i})$ to $\ell(i)$ in some expressions. Having selected stratum $\ell(i)$, we draw $\bm{x}_{i}\in^{d}$ via $\bm{x}_{i}=\phi_{\ell(i)}(\bm{u}_{i})$.
We now estimate $\mu={\mathbb{E}}(g(\bm{x}))$ by
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\omega(v_{i})g(\bm{x}_{i})\quad\text{where}\quad\omega(v_{i})=\frac{\alpha_{\ell(i)}}{\beta_{\ell(i)}},$ (2)
is an importance sampling weight. This weight only depends on the stratum $\ell$ and so we use $\omega_{\ell}$ for $\omega(v)$ when $\bm{\ell}(v)=\ell$. Two other ways to write this estimate will be useful. Because both $v_{i}$ and $\bm{x}_{i}$ are functions of $\bm{z}_{i}$ we write
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}f(\bm{z}_{i}),$ (3)
for
$f(\bm{z}_{i})=\omega(v_{i})g(\bm{x}_{i})=\omega(v_{i})g(\phi_{\bm{\ell}(v_{i})}(\bm{u}_{i})).$
This $f$ is our $s+1$ dimensional integrand for which we will use RQMC. It satisfies
$\mu=\int_{[0,1]^{s+1}}f(\bm{z})\mathrm{\,d}\bm{z}.$
We also define the within-stratum integrands individually as functions of $[0,1]^{s}$. The within-stratum integrands and integrals are
$h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))\quad\text{and}\quad\mu_{\ell}=\int_{[0,1]^{s}}h_{\ell}(\bm{u})\mathrm{\,d}\bm{u}$
respectively. We let
$S_{\ell}=\{1\leqslant i\leqslant n\mid\bm{\ell}(v_{i})=\ell\}$ (4)
index the set of sample points for stratum $\ell$ and define $n_{\ell}$ to be its cardinality. Now we can write
$\hat{\mu}$ $=\frac{1}{n}\sum_{i=1}^{n}\omega(v_{i})g(\bm{x}_{i})=\frac{1}{n}\sum_{i=1}^{n}\sum_{\ell=1}^{L}\bm{1}\{\bm{\ell}(v_{i})=\ell\}\omega_{\ell}h_{\ell}(\bm{u}_{i}).$
When every $n_{\ell}\geqslant 1$ we can write
$\hat{\mu}=\sum_{\ell=1}^{L}\omega_{\ell}\frac{n_{\ell}}{n}\hat{\mu}_{\ell}\quad\text{where}\quad\hat{\mu}_{\ell}=\frac{1}{n_{\ell}}\sum_{i\in S_{\ell}}h_{\ell}(\bm{u}_{i})$ (5)
is a within-stratum sample mean. For the RQMC method we consider we show below that $n_{\ell}\geqslant 1$ whenever $\beta_{\ell}>2/n$. We do not need $\Pr(n_{\ell}>0)=1$ for unbiasedness of $\hat{\mu}$, but without it the stratum sample means are not always well defined.
At this point we can introduce a small generalization. Since we already have a dependence on $\ell$ in $h_{\ell}$ we could also replace $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))$ by $h_{\ell}(\bm{u})=g_{\ell}(\phi_{\ell}(\bm{u}))$ to allow the function being applied to $\bm{x}$ to vary with $\ell$.
2.2 Sampling the strata
The way we sample strata does not affect MC computations but it makes a difference in RQMC. Having chosen $\beta_{\ell}>0$ with $\sum_{\ell=1}^{L}\beta_{\ell}=1$, we let $B_{0}=0$ and $B_{\ell}=\sum_{1\leqslant k\leqslant\ell}\beta_{k}$ for $1\leqslant\ell\leqslant L$. Then given $v$, we select stratum $\ell$ when
$B_{\ell-1}\leqslant v<B_{\ell}.$
While $v=1$ has probability zero, for completeness we set $\bm{\ell}(v)=L$ when $v=1$. Because $v\sim\mathbf{U}[0,1]$, the expected number of samples from stratum $\ell$ is $n\beta_{\ell}$. We will always order the strata so that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$. That brings an advantage in Section 2.5.
We will assume that the sampling values $v_{1},\dots,v_{n}$ that we use are ‘stratified’. By this we mean that every interval $I_{k}=[(k-1)/n,k/n)$ contains precisely one of $v_{1},\dots,v_{n}$ for $k=1,\dots,n$. The next proposition shows that if every $\beta_{\ell}>2/n$ then every $n_{\ell}>0$ when $v_{i}$ are stratified.
Proposition 1.
For $n\geqslant 1$, let $v_{1},\dots,v_{n}$ be stratified. For $[A,B)\subset[0,1)$, let $n_{*}$ be the number of $v_{i}\in[A,B)$. Then
$\lceil n\beta\rceil-2\leqslant n_{*}\leqslant\lfloor n\beta\rfloor+2$ (6)
where $\beta=B-A$.
Proof.
Write $A=r/n+a$ and $B=s/n+b$ for integers $0\leqslant r\leqslant s<n$ and $a,b\in[0,1/n)$. Then $\beta=t/n+c$ for $t=r-s$ and $-1/n<c<1/n$. If $r=s$, then $0\leqslant n_{*}\leqslant 1$ and then (6) holds trivially.
For $s>r$ we may write
$[A,B)=[A,(r+1)/n)\,\cup\,[(r+1)/n,s/n)\,\cup\,[s/n,s/n+b).$
Then $[A,B)$ contains at least $s-r-1$ of the intervals $I_{k}=[(k-1)/n,k/n)$ for $k=1,\dots,n$ and it intersects with at most $s-r+1$ of them. Therefore $|n_{*}-(s-r)|\leqslant 1$. Now $\lceil n\beta\rceil=s-r+\lceil nc\rceil$ and so $s-r\geqslant\lceil n\beta\rceil-1$ establishing the lower limit in (6). Similarly $\lfloor n\beta\rfloor=s-r+\lfloor nc\rfloor$ so $s-r\leqslant\lfloor n\beta\rfloor+1$ establishing the upper limit in (6). ∎
2.3 Local and star discrepancy
For $\bm{a}\in[0,1)^{d}$, the anchored box $[\bm{0},\bm{a})$ has volume $\mathrm{Vol}([\bm{0},\bm{a}))=\prod_{j=1}^{d}a_{j}$. The local discrepancy of $n$ points $\bm{x}_{1},\dots,\bm{x}_{n}\in[0,1)^{d}$ at $\bm{a}$ is
$\delta(\bm{a})=\frac{1}{n}\sum_{i=1}^{n}\bm{1}\{\bm{x}_{i}\in[\bm{0},\bm{a})\}-\mathrm{Vol}([\bm{0},\bm{a})).$
The star discrepancy of $\bm{x}_{1},\dots,\bm{x}_{n}$ is
$D_{n}^{*}=D_{n}^{*}(\bm{x}_{1},\dots,\bm{x}_{n})=\sup_{\bm{a}\in[0,1)^{d}}|\delta(\bm{a})|.$
The Koksma-Hlawka inequality (Hickernell, 2014) is that
$\Bigl{|}\frac{1}{n}\sum_{i=1}^{n}f(\bm{x}_{i})-\int_{[0,1]^{d}}f(\bm{x})\mathrm{\,d}\bm{x}\Bigr{|}\leqslant D_{n}^{*}(\bm{x}_{1},\dots,\bm{x}_{n})V_{{\mathrm{HK}}}(f)$ (7)
where $V_{{\mathrm{HK}}}$ is the total variation of $f$ in the sense of Hardy and Krause.
2.4 Digital nets
Here we present the defining properties of digital nets. They exist in integer bases $b\geqslant 2$ (Dick and Pillichshammer, 2010) but we only consider base $2$ here. For a dimension $d\geqslant 1$, elementary intervals in base $2$ are dyadic hyperrectangles of the form
$H(\bm{c},\bm{k})=\prod_{j=1}^{d}\Bigl{[}\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr{)}$
for integers $k_{j}\geqslant 0$ and $0\leqslant c_{j}<2^{k_{j}}$. We will use $|\bm{k}|=\sum_{j=1}^{d}k_{j}$, so $H(\bm{c},\bm{k})$ has volume $2^{-|\bm{k}|}$.
For integers $m\geqslant t\geqslant 0$ the $n=2^{m}$ points $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1)^{d}$ are a $(t,m,d)$-net in base $b$ if $H(\bm{c},\bm{k})$ contains precisely $nb^{-|\bm{k}|}=b^{m-|\bm{k}|}$ of them, whenever $m-|\bm{k}|\geqslant t$. Smaller $t$ are better. We always have $t\leqslant m$ because $t=m$ just means that all the points are in $[0,1)^{d}$.
A $(t,d)$-sequence in base $2$ is an infinite sequence of values $\bm{x}_{i}\in[0,1]^{d}$, with the following property: any subsequence $\bm{x}_{a2^{r}+1},\dots,\bm{x}_{(a+1)2^{r}}$ for integers $a\geqslant 0$ and $r\geqslant m$ is a $(t,r,d)$-sequence in base $2$. With a $(t,d)$-sequence we get an infinite sequence of $(t,m,d)$-nets. Consecutive pairs of those $(t,m,d)$-nets yield an infinite sequence of $(t,m+1,d)$-nets of which pairs form $(t,m+2,d)$-nets and so on without limit. We will use digital nets in base $2$ taken from Sobol’ sequences using the direction numbers from Joe and Kuo (2008).
We call a $(t,m,d)$-net in base $2$ stratified if for each component $j=1,\dots,d$, the values $x_{1j},\dots,x_{nj}$ are stratified. A net with $t=0$ is automatically stratified, but some digital nets with $t>0$ are also stratified. The $(t,m,d)$-nets from Sobol’ sequences are stratified. The stratification property means we could use the $j$’th variable in our $(t,m,d)$-net to select $\ell(i)$ for any $j$. Stratification implies the ‘projection regularity’ property of Sloan and Joe (1994) in which we never get $x_{ij}=x_{i^{\prime}j}$ for any $i\neq i^{\prime}$ and any $j$.
The Sobol’ points we use are ones that have been scrambled. The scrambles we consider are either nested uniform scrambles (Owen, 1995) or random linear scrambles with digital shifts (Matoušek, 1998). Scrambling makes each $\bm{z}_{i}\sim\mathbf{U}[0,1]^{d}$ individually while $\bm{z}_{1},\dots,\bm{z}_{n}$ remain a $(t,m,d)$-net collectively.
We will estimate $\mu$ by $\hat{\mu}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are a scrambled $(t,m,s+1)$-net in base $2$. This scrambled net integration has several desirable properties listed in Owen and Rudolf (2020):
  1. 1)
    If $f\in L_{1}[0,1]^{s+1}$, then ${\mathbb{E}}(\hat{\mu})=\mu$.
  2. 2)
    If $f\in L_{1+\epsilon}[0,1]^{s+1}$ for some $\epsilon>0$ , then $\Pr(\lim_{n\to\infty}\hat{\mu}=\mu)=1$.
  3. 3)
    If $f\in L_{2}[0,1]^{s+1}$, then ${\mathrm{Var}}(\hat{\mu})=o(1/n)$ as $n\to\infty$.
  4. 4)
    When every partial derivative of $f$ taken up to one time with respect to each component of $\bm{z}$ is in $L_{2}$ then ${\mathrm{Var}}(\hat{\mu})=O(n^{-3+\epsilon})$.
  5. 5)
    There exists a finite $\Gamma$ with ${\mathrm{Var}}(\hat{\mu})\leqslant\Gamma\sigma^{2}/n$ where $\sigma^{2}/n$ is the variance under ordinary Monte Carlo (MC) sampling.
The scrambled Sobol’ nets are stratified.
From point 3 above we know that ${\mathrm{Var}}(\hat{\mu})=o(n^{-1})$ for $f\in L_{2}$. The Sobol’ nets have $D_{n}^{*}=O(n^{-1}\log(n)^{d-1})$ (Niederreiter, 1992, Theorem 4.10). Scrambled nets are still nets, so this holds for them too. Then by the Koksma-Hlawka inequality, ${\mathrm{Var}}(\hat{\mu})=O(n^{-2+\epsilon})$ holds for any $\epsilon>0$ when $V_{{\mathrm{HK}}}(f)<\infty$, which we describe as $f$ being in BVHK. For regular $f$ described in point 4 above we have ${\mathrm{Var}}(\hat{\mu})=O(n^{-3+\epsilon})$. Higher order digital nets can attain even better convergence rates under correspondingly stronger regularity (Dick, 2011), though they are not widely used for higher dimensions. It is common to write these powers of $\log(n)$ as $O(n^{\epsilon})$. For a discussion of that point, see Owen and Pan (2022).
2.5 Discrepancy within each stratum
Here we consider how uniformly distributed the points in each stratum are. The first case we consider has $\beta_{\ell}$ decreasing with $\ell$ and equal to negative powers of $2$. We show how to find such $\beta_{\ell}$ in Section 4. Then the sample points within any given stratum are scrambled nets, and the error can be as good as $O(n^{-3+\epsilon})$ for regular integrands. After that, we consider arbitrary $\beta_{\ell}>2/n$. Then the points in a given stratum still have small discrepancy but they are not generally scrambled digital nets. In that case the rate in stratum $\ell$ is $O(n_{\ell}^{-2+\epsilon})$ for within-stratum integrands that are in BVHK.
Proposition 2.
For $n=2^{m}$ and $s\geqslant 1$, let $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1]^{s+1}$ be a stratified $(t,m,s+1)$-net in base $2$. For $L\geqslant 2$, let $\beta_{\ell}=2^{-\kappa_{\ell}}$ for $1\leqslant\kappa_{1}\leqslant\kappa_{2}\leqslant\cdots\leqslant\kappa_{L}<m$ satisfy $\sum_{\ell=1}^{L}\beta_{\ell}=1$. Write $\bm{z}_{i}=(v_{i},u_{i1},\dots,u_{is})$ for $\bm{u}_{i}\in[0,1]^{s}$. Then for each $\ell=1,\dots,L$, the points $\bm{u}_{i}$ for which $\bm{\ell}(v_{i})=\ell$ form a $(t_{\ell},m_{\ell},s)$-net in base $2$ with $m_{\ell}=m-\kappa_{\ell}$ and $t_{\ell}=\min(t,m_{\ell})$.
Proof.
For $S_{\ell}$ defined at (4), sample $i\in S_{\ell}$ holds if and only if $B_{\ell-1}\leqslant z_{i1}<B_{\ell}$ or equivalently
$\sum_{1\leqslant r<\ell}2^{-\kappa_{r}}\leqslant z_{i1}<\sum_{1\leqslant r\leqslant\ell}2^{-\kappa_{r}}$
which we can write as
$\frac{\gamma}{2^{\kappa_{\ell}}}\leqslant z_{i1}<\frac{\gamma+1}{2^{\kappa_{\ell}}}$ (8)
for integer $\gamma=\sum_{1\leqslant r<\ell}2^{\kappa_{\ell}-\kappa_{r}}\geqslant 0$. Because the net is stratified, the number of observations $i$ that satisfy equation (8) is $2^{m-\kappa_{\ell}}=2^{m_{\ell}}$. Then $\{u_{i}\mid i\in S_{\ell}\}\subset[0,1)^{s}$ has cardinality $2^{m_{\ell}}$ making it a $(t_{\ell},m_{\ell},s)$-net in base $2$ for some value $t_{\ell}\leqslant m_{\ell}$.
Now consider a dyadic box in $[0,1]^{s}$ of the form
$\prod_{j=1}^{s}\Bigl{[}\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr{)}.$
This box has volume $2^{-|\bm{k}|}$. The number of points $\bm{u}_{i}$ from stratum $\ell$ in this box equals the number of points $\bm{z}_{i}\in[0,1)^{s+1}$ that belong to
$\Bigl{[}\frac{\gamma}{2^{\kappa_{\ell}}},\frac{\gamma+1}{2^{\kappa_{\ell}}}\Bigr{)}\times\prod_{j=1}^{s}\Bigl{[}\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr{)}.$
If $m-\kappa_{\ell}-|\bm{k}|\geqslant t$, then by the digital net property of $\bm{z}_{1},\dots,\bm{z}_{n}$, this number is $2^{m-\kappa_{\ell}-|\bm{k}|}$. That is, if $m_{\ell}\geqslant|\bm{k}|+t$, then the number of such points is $2^{m_{\ell}-|\bm{k}|}$
To motivate the constraint that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$, consider what would happen if we took $\beta_{1}=1/8$, $\beta_{2}=1/2$, $\beta_{3}=1/4$ and $\beta_{4}=1/8$. Then the points for the second stratum would arise from $v\in[1/8,5/8)$. This is not an elementary interval in base $2$, but is instead the union of four such sets each of length $1/8$. Using the proper ordering $\beta_{1}=1/2$, $\beta_{2}=1/4$, $\beta_{3}=1/8$ and $\beta_{4}=1/8$, we get intervals $[0,1/2)$, $[1/2,3/4)$, $[3/4,7/8)$ and $[7/8,1)$ for the four strata and these are all elementary intervals in base $2$. The digital net property assures a better equidistribution property for an elementary interval than it does for an equal sized union of two or more elementary intervals.
Now we consider what happens when $\beta_{\ell}$ are not negative powers of two. We get low discrepancy within strata but we do not get digital nets there.
Proposition 3.
Let $\beta_{1},\dots,\beta_{L}\in(0,1)$ sum to one. Let $\bm{z}_{1},\dots,\bm{z}_{n}$ be a stratified $(t,m,s+1)$-net in base $2$ with $n>2/\min_{\ell}\beta_{\ell}$. Then $n_{\ell}>0$ and the star discrepancy of $\{\bm{u}_{i}\mid i\in S_{\ell}\}$ is $O(n_{\ell}^{-1}\log(n_{\ell})^{s})$.
Proof.
The local discrepancy of $\bm{z}_{1},\dots,\bm{z}_{n}$ at the point $(b,\bm{c})\in[0,1]^{s+1}$ is
$\delta(b,\bm{c})=\frac{1}{n}\sum_{i=1}^{n}\bm{1}\{0\leqslant v_{i}<b\}\bm{1}\{\bm{z}_{i}\in[\bm{0},\bm{c})\}-b\times\mathrm{Vol}([\bm{0},\bm{c}))$
where $\mathrm{Vol}([\bm{0},\bm{c}))=\prod_{j=1}^{s}c_{j}$. Because $\bm{z}_{1},\dots,\bm{z}_{n}$ are a digital net, $|\delta(b,\bm{c})|=O(n^{-1}\log(n)^{s})$.
The sample points for stratum $\ell$ are the $n_{\ell}$ points $\bm{x}_{i}$ for which $B_{\ell-1}\leqslant v_{i}<B_{\ell}$. By equation (6), $n_{\ell}=n\beta_{\ell}+O(1)$. The local discrepancy within stratum $\ell$ at $\bm{c}\in[0,1)^{s}$ is
$\ \quad\frac{1}{n_{\ell}}\sum_{i=1}^{n}\bm{1}\{B_{\ell-1}\leqslant v_{i}<B_{\ell}\}\bm{1}\{\bm{x}_{i}\in[\bm{0},\bm{c})\}-\mathrm{Vol}([\bm{0},\bm{c}))$
$=\frac{n}{n_{\ell}}\Bigl{(}\beta_{\ell}\mathrm{Vol}([\bm{0},\bm{c}))+\delta(B_{\ell},\bm{c})-\delta(B_{\ell-1},\bm{c})\Bigr{)}-\mathrm{Vol}([\bm{0},\bm{c})).$
Its absolute value is at most
$\Bigl{|}\frac{n\beta_{\ell}}{n}-1\Bigr{|}\mathrm{Vol}([\bm{0},\bm{c}))+\frac{2n}{n_{\ell}}O(n^{-1}\log(n)^{s})$ $=O(n_{\ell}^{-1}\log(n_{\ell}/\beta_{\ell})^{s})$
$=O(n_{\ell}^{-1}\log(n_{\ell})^{s})$
and so the star discrepancy of $\bm{u}_{i}$ for $i\in S_{\ell}$ is $O(n_{\ell}^{-1}\log(n_{\ell})^{s})$. ∎
If $\bm{z}_{1},\dots,\bm{z}_{n}$ are Halton points (scrambled or otherwise) with $n=2^{m}$ then $\bm{z}_{i1}$ for $1\leqslant i\leqslant n$ are stratified.
3 Choosing $\beta_{\ell}$ for given $\alpha_{\ell}$
Here we develop some principled criteria for choosing $\beta_{\ell}$. Our working model is that
${\mathrm{Var}}(\hat{\mu}_{\ell})=\tau_{\ell}n_{\ell}^{-\rho}$ (9)
for some $\rho>0$ and $\tau_{\ell}>0$. For simplicity, we assume that every stratum has the same rate $\rho$, apart from a remark at the end of Section 3.1. The appropriate value of $\rho$ is not necessarily known at the time we choose $\beta_{\ell}$. We will show that choosing $\beta_{\ell}$ based on the wrong value of $\rho$ can adversely affect the implied constant in the asymptotic MSE while leaving the convergence rate unchanged.
Values of $\rho$ equal to $1$, $2$ and $3$ correspond to upper bounds on the covergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases. It often happens that the empirical MSE of an RQMC rule follows an apparent rate of $n^{-\rho}$ for a value of $\rho$ that is not close to an integer. That can happen when the practically usable values of $n$ are smaller than the sample size at which the asymptotic result holds. Unbounded integrands are not in BVHK, and yet they will have an MSE that is $o(n^{-1})$. Those often appear to have $21<\rho<$ 2.
Our sampling algorithm will deliver $n_{\ell}$ with $|n_{\ell}-n\beta_{\ell}|\leqslant 2$ by Proposition 1. For $\rho<2$, we will work as if $n_{\ell}=n\beta_{\ell}$, exactly. Changing one or two points per stratum will make a difference of $O(1/n)$ to the integral estimate for a bounded integrand and will have a root mean squared difference of this magnitude for finite variances. The overall root mean squared error (RMSE) is $O(n^{-\rho/2})$ and so for $\rho<2$ these small differences in $n_{\ell}$ are negligible.
For $\rho>2$, we have the reverse situation. Changing $n_{\ell}$ by even ±$\pm 1$ can worsen the convergence rate because changing $\hat{\mu}$ by $O(1/n)$ is a bigger change than the RMSE of $O(n^{-\rho/2})$. A convergence rate with $\rho>2$ for scrambled Sobol’ points is only available for sample sizes that are powers of two. Section 4 shows how to obtain $n_{\ell}$ equal to powers of two.
With $n_{\ell}=n\beta_{\ell}$ and $\omega_{\ell}=\alpha_{\ell}/\beta_{\ell}$, we can rewrite (5) as
$\hat{\mu}=\sum_{\ell=1}^{L}\omega_{\ell}\frac{n_{\ell}}{n}\hat{\mu}_{\ell}=\sum_{\ell=1}^{L}\alpha_{\ell}\hat{\mu}_{\ell}.$
The importance sampling ratio $\omega_{\ell}$ has canceled from the expression for $\hat{\mu}$, leaving a simply weighted some of within-stratum means.
3.1 Optimal sample sizes
Under the model (9),
${\mathrm{Var}}(\hat{\mu})={\mathrm{Var}}\biggl{(}\sum_{\ell=1}^{L}\alpha_{\ell}\hat{\mu}_{\ell}\biggr{)}\leqslant\biggl{(}\sum_{\ell=1}^{L}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}\biggr{)}^{2}$ (10)
where the upper bound is from the very conservative assumption that the estimates within all $L$ strata have unit correlations. With no prior information on the relative sizes of $\tau_{\ell}$ we adopt
$C_{1}(\rho)=C_{1}(\rho;n_{1},\dots,n_{L})=\sum_{\ell=1}^{L}\alpha_{\ell}n_{\ell}^{-\rho/2}$ (11)
as a criterion to minimize subject to $\sum_{\ell=1}^{L}n_{\ell}=n$. We form the Lagrangian
$G_{1}=\sum_{\ell=1}^{L}\alpha_{\ell}n_{\ell}^{-\rho/2}+\lambda\sum_{\ell=1}^{L}n_{\ell}$
and setting $\partial G_{1}/\partial n_{\ell}$ to zero yields a solution with
$n_{\ell}\propto\alpha_{\ell}^{2/(\rho+2)}.$ (12)
We then suppose that the user chooses integers $n_{\ell}$ summing to $n=2^{m}$ that are nearly proportional to $\alpha_{\ell}^{2/(\rho+2)}$.
RQMC points have a space filling property. We expect that the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$. We might then anticipate primarily non-positive correlations among the $\hat{\mu}_{\ell}$. This motivates a different criterion
$C_{0}(\rho)=C_{0}(\rho;n_{1},\dots,n_{L})=\sum_{\ell=1}^{L}\alpha_{\ell}^{2}n_{\ell}^{-\rho}$ (13)
based on zero correlations among the $\hat{\mu}_{\ell}$. Forming the Lagrangian for this criterion and setting to zero the partial derivatives with respect to $n_{\ell}$ leads to
$n_{\ell}\propto\alpha_{\ell}^{2/(\rho+1)}.$ (14)
Higher convergence rates $\rho$ tend to equalize the $n_{\ell}$. We can think of $n_{\ell}=n/L$ as arising in an $\rho\to\infty$ limit. We will also see equal $n_{\ell}$ in a minimax argument in Section 5.
We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for ansatzs 00 and $1$, and three different hypotheses on $\rho$. They lead to the four different rates given in Table 1. The case for ansatz 00 and $\rho=1$ (uncorrelated estimates at the MC rate) is the only one with a proportional allocation. The other choices all allocate more than a proportional number of samples to the strata with smaller $\alpha_{\ell}$. This is even true for the plain MC rate ($\rho=1$) under the conservative ansatz 1.
When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then equations (12) and (14) don’t distinguish betweem different values of $\rho$. After raising $\alpha_{\ell}$ to any power they are still equal and then a near equal integer allocation is appropriate.
When we do have information about the relative sizes of $\tau_{\ell}$, then we can replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}$ in the formulas in Table 1. In that case, the proportionality for $C_{0}$ and $\rho=1$ matches the Neyman allocation from survey sampling, in Neyman (1934). When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$ then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}(\tau_{\ell}/c_{\ell})^{1/2}$ in Table 1.
$\rho:$ 1 2 3
$C_{0}$ $\alpha_{\ell}$ $\alpha_{\ell}^{2/3}$ $\alpha_{\ell}^{1/2}$
$C_{1}$ $\alpha_{\ell}^{2/3}$ $\alpha_{\ell}^{1/2}$ $\alpha_{\ell}^{2/5}$
Table 1: This table shows ideal rates for $n_{\ell}$ as functions of $\alpha_{\ell}$. Under ansatz 00, estimates from different strata are uncorrelated. Ansatz $1$ uses a bound assuming that those estimates are perfectly correlated. These rates apply when the variance within stratum $\ell$ decreases proportionally to $n_{\ell}^{-\rho}$.
If $\rho$ varies from stratum, taking the value $\rho_{\ell}$ in stratum $\ell$, then the Lagrange multiplier argument leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho_{\ell}+1)}$. Strata with higher $\rho_{\ell}$ get larger $n_{\ell}$ and $\beta_{\ell}=n_{\ell}/n$ does not depend on $n$.
3.2 Suboptimality
Now suppose that ${\mathrm{Var}}(\hat{\mu}_{\ell})\propto n_{\ell}^{-\rho}$ but due to imperfect knowledge, we work as if ${\mathrm{Var}}(\hat{\mu}_{\ell})\propto n_{\ell}^{-\gamma}$. We measure inefficiency by the ratio of the MSE we get if we design with a value $\gamma$ when we should have used $\rho$, so smaller values are better, and $1$ is best possible.
In ansatz 0 we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
$I_{0}(\gamma\!\mid\!\rho)$ $=\frac{\sum_{\ell=1}^{L}\alpha^{2}_{\ell}\Bigl{(}n\alpha_{\ell}^{2/(\gamma+1)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+1)}\Bigr{)}^{-\rho}}{\sum_{\ell=1}^{L}\alpha^{2}_{\ell}\Bigl{(}n\alpha_{\ell}^{2/(\rho+1)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)}\Bigr{)}^{-\rho}}$
$=\sum_{\ell=1}^{L}\alpha_{\ell}^{2-2\rho/(\gamma+1)}\frac{\bigl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+1)}\bigr{)}^{\rho}}{\bigl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)}\bigr{)}^{\rho+1}}.$ (15)
The factor $n^{-\rho}$ has canceled from the numerator and denominator and so we get the same rate of convergence with an incorrect $\gamma$ as we would with the optimal $\gamma=\rho$. Only the lead constant is affected.
For ansatz 1, assuming a rate $n_{\ell}^{-\gamma}$ instead of $n_{\ell}^{-\rho}$ would lead to $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+2)}$. To put the inefficiency in squared error terms, comparable to $I_{0}$ we use $C^{2}_{1}(\gamma)/C^{2}_{1}(\rho)$, getting
$I_{1}(\gamma\!\mid\!\rho)$ $=\left(\frac{\sum_{\ell=1}^{L}\alpha_{\ell}\bigl{(}n\alpha_{\ell}^{2/(\gamma+2)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+2)}\bigr{)}^{-\rho/2}}{\sum_{\ell=1}^{L}\alpha_{\ell}\bigl{(}n\alpha_{\ell}^{2/(\rho+2)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}\bigr{)}^{-\rho/2}}\right)^{2}$
$=\frac{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{1-\rho/(\gamma+2)}\bigr{)}^{2}\bigl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+2)}\bigr{)}^{\rho}}{\bigl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}\bigr{)}^{\rho+2}}.$
We think that ansatz 0 makes a more reasonable assumption than ansatz 1. Then a cautious choice is to take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma_{0}+1)}$ where
$\gamma_{0}=\operatorname*{argmin}_{1\leqslant\gamma\leqslant 3}\max_{1\leqslant\rho\leqslant 3}I_{0}(\gamma\!\mid\!\rho).$ (16)
When we know that $\rho\leqslant 2$ we might only let $\rho$ and $\gamma$ belong to $[1,2]$.
The minimax approach of equations (16) gives us ways to choose $\gamma$ for a given list of values $\alpha_{1},\dots,\alpha_{L}$. It often happens that the worst value of $\rho$ for given $\gamma$ is one of the allowed extremes, but this does not always happen. We have seen such exceptions when one of the $\alpha_{\ell}$ is much larger than the others. In our numerical examples the maximum over $\rho$ of $I_{0}(2\!\mid\!\rho)$ or $I_{0}(3\!\mid\!\rho)$ was not much larger than $I_{0}(\gamma_{0}\!\mid\!\rho)$.
Next we consider recommendations that can be made generally without specific values of $\alpha_{\ell}$ in mind. The first thing to notice is that both $I_{0}$ and $I_{1}$ contain a numerator sum that could potentially have a negative power of $\alpha_{\ell}$. For general use where some $\alpha_{\ell}$ might be small, we want to avoid that possibility. There is no possibility of an even more negative power in the denominator to compensate. To keep a positive exponent in the numerator of $I_{0}$ we must have $2-2\rho/(\gamma+1)\geqslant 0$. After rearrangement this becomes a constraint that $\gamma\geqslant\rho-1$, whatever the unknown $\rho$ might be. When $\rho\doteq 3$ is a possibility then we should take $\gamma\geqslant 2$. If instead we use $\gamma=1$ in a $\rho=3$ setting then the first factor in $I_{0}$ is greater than $\alpha_{L}^{-1}$. For $I_{1}$ we want $1-\rho/(\gamma+2)\geqslant 0$ or $\gamma\geqslant\rho-2$. Since we already had $\gamma\geqslant\rho-1$ using ansatz 0, this is not a further constraint. When we know that $\rho\leqslant 2$, then any $\gamma\geqslant 1$ avoids a negative exponent for $\alpha_{\ell}$.
When $\rho=1$ (MC rate), then
$I_{0}(2\!\mid\!1)$ $=\sum_{\ell=1}^{L}\alpha_{\ell}^{4/3}\sum_{\ell=1}^{L}\alpha_{\ell}^{2/3},\quad\text{and}$
$I_{0}(3\!\mid\!1)$ $=\sum_{\ell=1}^{L}\alpha_{\ell}^{3/2}\sum_{\ell=1}^{L}\alpha_{\ell}^{1/2}.$
It is not difficult to show that $I_{0}(3\!\mid\!1)\geqslant I_{0}(2\!\mid\!1)\geqslant I_{0}(1\!\mid\!1)=1$. Proposition 4 shows that more general monotonicities hold.
Proposition 4.
For $L>1$, let $\alpha_{\ell}>0$ for $\ell=1,\dots,L$ sum to one. If $\rho\geqslant 1$ then
$I_{0}(\gamma_{1}\!\mid\!\rho)\geqslant I_{0}(\gamma_{2}\!\mid\!\rho)\geqslant I_{0}(\rho\!\mid\!\rho)$ (17)
whenever either $\gamma_{1}\geqslant\gamma_{2}\geqslant\rho$ or $1\leqslant\gamma_{1}\leqslant\gamma_{2}\leqslant\rho$.
Proof.
See Appendix A. ∎
If the unknown true rate is $\rho=2$, we are better off choosing $\gamma=3$ than $\gamma=1$. Proposition 5 has this ‘bet on optimism’ result.
Proposition 5.
For any $\alpha_{\ell}>0$ with $\sum_{\ell=1}^{L}\alpha_{\ell}=1$,
$I_{0}(3\!\mid\!2)\leqslant I_{0}(1\!\mid\!2).$
Proof.
From equation (3.2)
$I_{0}(1\!\mid\!2)=\frac{L}{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{2/3}\bigr{)}^{3}}\quad\text{and}\quad I_{0}(3\!\mid\!2)=\frac{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{1/2}\bigr{)}^{2}}{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{2/3}\bigr{)}^{3}}.$
These expressions have the same denominator and then the result follows from the Cauchy-Schwarz inequality applied to the numerators. ∎
4 Making each $n_{\ell}$ a power of two
Here we consider how to allocate sample sizes $n_{\ell}=2^{m_{\ell}}$ that sum to given a total sample size $n=2^{m}$. We assume throughout that $n\geqslant L$ and that we want every $n_{\ell}\geqslant 1$. The advantage of having every $n_{\ell}$ be a power of two is that we then get scrambled net sampling within each of the $L$ strata as shown by Proposition 2. Sample sizes that are a power of two are a best practice in scrambled Sobol’ sampling (Owen, 2020). The results with MSE $O(n^{-3+\epsilon})$ for scrambled Sobol’ points require such sample sizes.
The potential disadvantage of having $n_{\ell}$ be a power of $2$ is that we have many fewer sampling ratios $\alpha_{\ell}/\beta_{\ell}$ at our disposal. For example, with $L=2$, we would have to use $n_{1}=n_{2}=n/2$ and this could be a bad fit for $\alpha_{1}=0.9999$ and $\alpha_{2}=0.0001$.
We take $\beta_{\ell}=2^{-\kappa_{\ell}}$ such that $\sum_{\ell=1}^{L}\beta_{\ell}=1$. That is, $\beta_{\ell}$ are a partition of $1$. Partitions like $(1/2,1/4,1/4)$ and $(1/4,1/2,1/4)$ are equivalent because any assignment of one of those to $\beta_{\ell}$ can be achieved by the other after a permutation of indices.
We suppose that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$ for $L\geqslant 2$ and look at sample fractions $(n_{1}/n,n_{2}/n,\dots,n_{L}/n)$. For $L=2$ the only solution is $(1/2,1/2)$ and for $L=3$ the only solution is $(1/2,1/4,1/4)$. For $L=4$, there are two possibilities: $(1/2,1/4,1/8,1/8)$ and $(1/4,1/4,1/4,1/4)$ and for $5\leqslant L\leqslant 10$ the numbers of possibilities are: 3, 5, 9, 16, and 28. The tables in Figure 1 show the partitions for $4\leqslant L\leqslant 8$.
L Partitions: 4 $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ 5 $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{2}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ 6 $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{2}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ 7 $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{64}$ $\tfrac{1}{64}$ $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{2}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{2}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$             L Partitions: 8 $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{64}$ $\tfrac{1}{128}$ $\tfrac{1}{128}$ $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{64}$ $\tfrac{1}{64}$ $\tfrac{1}{64}$ $\tfrac{1}{64}$ $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{64}$ $\tfrac{1}{64}$ $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{64}$ $\tfrac{1}{64}$ $\tfrac{1}{2}$ $\tfrac{1}{4}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{2}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{64}$ $\tfrac{1}{64}$ $\tfrac{1}{2}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{2}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{2}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{64}$ $\tfrac{1}{64}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{32}$ $\tfrac{1}{32}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{4}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{16}$ $\tfrac{1}{16}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$ $\tfrac{1}{8}$
Figure 1: These are the partitions of unity into $L$ negative powers of $2$ for $L\in\{4,5,6,7,8\}$.
The number of such partitions as a function of $L$ is given by sequence A002572 at the online encyclopedia of integer sequences, (OEIS Foundation Inc., 2023) along with motivating applications such as counting the number of rooted binary trees or, more recently, counting nonequivalent compact Huffman codes (Elsholtz et al. (2023)). The latter reference mentions that the number of solutions is asymptotic to $R\theta^{L}$ for $R\doteq 0.14$ and $\theta\doteq 1.794$. This is a somewhat faster growth rate than the Fibonacci numbers have: their counterpart to $\theta$ is $(1+\sqrt{5})/2\doteq 1.618$. That asymptote is not a good description for small $L$. For $L=10$, it is over 4000, while the exact value is 28.
To choose $\beta_{\ell}$, or equivalently $n_{\ell}$, we uses a forward stepwise algorithm. We initialize every $m_{\ell}=0$ so every $n_{\ell}=2^{m_{\ell}}$ is initially $1$. At every step of the algorithm there are $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$ samples left to allocate. If $n^{\prime}=0$, the algorithm terminates. If $n^{\prime}>0$ then we select one of the indices $\ell$ with $n_{\ell}\leqslant n^{\prime}$ and double that $n_{\ell}$, replacing $m_{\ell}$ by $m_{\ell}+1$. Before describing which $n_{\ell}$ to double, we show that any such algorithm always terminates.
Proposition 6.
The above algorithm always terminates in a finite number of steps.
Proof.
First, if $n^{\prime}>0$ then there is an ‘eligible’ index $\ell$ with $n_{\ell}\leqslant n^{\prime}$, that can be doubled without causing $\sum_{\ell=1}^{L}n_{\ell}>n$. At any stage in the algorithm, $n_{\ell^{*}}$ for $\ell^{*}=\operatorname*{argmin}_{1\leqslant\ell\leqslant L}n_{\ell}$ is a power of two. It divides $n$ and it divides every $n_{\ell}$, so it divides the nonzero value $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$, hence $n_{\ell^{*}}\leqslant n^{\prime}$. Finally, every step with $n^{\prime}>0$ reduces $n^{\prime}$, and so $n^{\prime}$ must reach zero in a finite number of steps starting from $n^{\prime}=n-L$. ∎
Under ansatz 0, the ideal proportion of samples in stratum $\ell$ is
$\xi_{\ell,\rho}=\alpha_{\ell}^{2/(\rho+1)}\Bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)}.$
We choose to double the sample size for stratum
$\ell^{*}_{0}=\operatorname*{argmax}_{1\leqslant\ell\leqslant L}\frac{\xi_{\ell,\rho}}{2^{m_{\ell}}}\bm{1}\Bigl{\{}2^{m_{\ell}}\leqslant n-\sum_{j=1}^{L}2^{m_{j}}\Bigr{\}}$
which we consider to be the most ‘under-represented’ of the eligible strata. The analogous algorithm under ansatz $1$ is to double the sample size of stratum
$\ell^{*}_{1}=\operatorname*{argmax}_{1\leqslant\ell\leqslant L}\frac{\xi_{\ell,\rho}}{2^{m_{\ell}}}\bm{1}\Bigl{\{}2^{m_{\ell}}\leqslant n-\sum_{j=1}^{L}2^{m_{j}}\Bigr{\}}$
where
$\xi_{\ell,\rho}=\alpha_{\ell}^{2/(\rho+2)}\Bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}.$
The pseudo-code for forward allocation algorithm for ansatz 0 is given in Algorithm 1.
Algorithm 1 Forward Stratified Allocation
0:  Budget $n=2^{m}\geqslant L$, stratum weights $\alpha_{1},\dots,\alpha_{L}$, parameter $\rho\geqslant 1$
0:  Powers $\bm{m}=(m_{1},\dots,m_{L})$ such that $n_{\ell}=2^{m_{\ell}}$ and $\sum_{\ell=1}^{L}n_{\ell}=n$
1:  Initialize $m_{\ell}\leftarrow 0$ for all $\ell=1,\dots,L$
2:  Compute $\xi_{\ell,\rho}\leftarrow\alpha_{\ell}^{2/(\rho+1)}\big{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)}$ for all $\ell$
3:  while $n-\sum_{\ell=1}^{L}2^{m_{\ell}}>0$ do
4:     Find $\ell^{*}\leftarrow\operatorname*{argmax}_{1\leqslant\ell\leqslant L}\left\{\dfrac{\xi_{\ell,\rho}}{2^{m_{\ell}}}\cdot\bm{1}\left\{2^{m_{\ell}}\leqslant n-\sum_{j=1}^{L}2^{m_{j}}\right\}\right\}$
5:     Update $m_{\ell^{*}}\leftarrow m_{\ell^{*}}+1$
6:  end while
7:  return  $\bm{m}=(m_{1},\dots,m_{L})$
5 Minimax choice
Here we show a minimax property of near-equal allocation. We do not assume equality among the $\tau_{\ell}$. For any integer $N\geqslant L$, let
$\Delta_{N}$ =$=\biggl{\{}\bm{n}=(n_{1},\dots,n_{L})\in\mathbb{N}^{L}\Bigm{|}\sum_{\ell=1}^{L}n_{\ell}=N\biggr{\}},\quad\text{and}\quad$
$\mathcal{T}$ =$=\Bigl{\{}\bm{\tau}=(\tau_{1},\dots,\tau_{L})\in[0,\infty)^{L}\bigm{|}\max_{1\leqslant\ell\leqslant L}\tau_{\ell}>0\Bigr{\}}$
be spaces of feasible sample sizes and possible variance constants, respectively. We allow some $\tau_{\ell}=0$ but there must be at least one $\tau_{\ell}>0$ to make the problem interesting.
We consider the increase in variance that we get using sample sizes $\bm{n}$ compared to some better sample sizes $\tilde{\bm{n}}$ that we might otherwise have used. For ansatzes $a=0,1$ we define suboptimality ratios
$R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}}{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}\tilde{n}_{\ell}^{-\rho/2}}\quad\text{and}\quad R_{0}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}n_{\ell}^{-\rho}}{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}\tilde{n}_{\ell}^{-\rho}}$
that compare different sample sizes taking account of the stratum variances $\tau_{\ell}$. These ratios depend on $\rho$ and $\alpha_{1},\dots,\alpha_{L}$, but we leave those out of the notation because they are held constant in this section.
Proposition 7.
For both ansatzes $a\in\{0,1\}$,
$\min_{\bm{n}\in\Delta_{N}}\max_{\tilde{\bm{n}}\in\Delta_{N}}\max_{\bm{\tau}\in\mathcal{T}}R_{a}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau})=R_{a}(\bm{n}^{*}\!\mid\!\tilde{\bm{n}}^{*};\bm{\tau}^{*})$ (18)
for minimax sample sizes
$n_{\ell}^{*}=\begin{cases}\left\lfloor{\frac{N}{L}}\right\rfloor+1,&\textrm{for }\ell=1,\dots,r\\ \left\lfloor{\frac{N}{L}}\right\rfloor,&\textrm{for }\ell=r+1,\dots,L\end{cases}$
with $r=N-L\lfloor N/L\rfloor$, worst case within-stratum variances $\tau_{\ell}=\lambda\bm{1}\{\ell=\ell_{*}$} for any $\lambda>0$ and worst case alternative sample sizes
$\begin{array}[]{ll}\tilde{n}_{\ell}^{*}=\begin{cases}N-L+1,&\text{if }\ell=\ell^{*}\\ 1,&\text{if }\ell\in\{1,2,\dots,L\}\setminus\{\ell^{*}\}\end{cases}\end{array}$
for any $\ell^{*}\in\{r+1,\dots,L\}$.
Proof.
The same argument applies to both ansatzes. We prove it for ansatz 1. Fixing $\bm{n}$ and $\tilde{\bm{n}}$ we see that the maximum over $\bm{\tau}$ of $R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau})$ arises when $\tau_{\ell}=0$ for all $\ell$ except $\ell^{*}(\bm{n},\tilde{\bm{n}})=\operatorname*{argmax}_{\ell}(\tilde{n}_{\ell}/n_{\ell})^{\rho/2}$ with $\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}$ then taking any value $\lambda>0$.
Now, we can go back to the initial problem and rewrite (18) as:
$\min_{\bm{n}\in\Delta_{N}}\max_{\tilde{\bm{n}}\in\Delta_{N}}\Bigl{(}\frac{\tilde{n}_{\ell^{*}(\bm{n},\tilde{\bm{n}})}}{{n}_{\ell^{*}(\bm{n},\tilde{\bm{n}})}}\Bigr{)}^{\rho/2}$
after noting that the factor $\alpha_{\ell^{*}(\bm{n},\tilde{\bm{n}})}\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}^{1/2}$ cancels from numerator and denominator. To maximize this ratio over $\tilde{\bm{n}}$ one takes $\tilde{\bm{n}}$ to be the vector with value $N-L+1$ in component $\ell^{*}=\operatorname*{argmin}_{\ell}n_{\ell}$ and ones elsewhere. Then the minimax problem becomes $\min_{\bm{n}\in\Delta_{N}}\bigl{(}(N-L+1)/\min_{1\leqslant\ell\leqslant N}n_{\ell}\bigr{)}^{\rho/2}$. The solution is to maximize the minimum value of $n_{\ell}$. We do that by making $\min_{\ell}n_{\ell}=\lfloor N/L\rfloor$. ∎
When we must have $n_{\ell}=2^{m_{\ell}}$ for integers $m_{\ell}\geqslant 0$, there is a corresponding minimax result that takes the $n_{\ell}$ as nearly equal as possible. Let $r=\lceil\log_{2}(L)\rceil$ and then every $\beta_{\ell}$ is either $2^{-r}$ or $2^{1-r}$. Solving $s2^{1-r}+(L-s)2^{-r}=1$, we see that there should be $s=2^{r}-L$ strata with the larger allocation. Then we take $n_{\ell}=n\beta_{\ell}$ for
$\beta_{\ell}=\begin{cases}2^{1-r},&1\leqslant\ell\leqslant 2^{r}-L\\ 2^{-r},&2^{r}-L<\ell\leqslant L.\end{cases}$
6 A toy example
Here we consider a very small example with only one continuous variable and $8$ strata for illustration. In stratum $\ell$, the parameter $\theta$ takes the value $\theta_{\ell}$ and then $x_{1}\sim\mathcal{N}(\theta_{\ell},1)$. This example uses
$\begin{split}\alpha&=(0.50,0.44,0.01,0.01,0.01,0.01,0.01,0.01),\quad\text{and}\\ \theta&=(0.7,1.0,1.5,1.6,1.7,1.8,1.9,2.0).\end{split}$ (19)
The within-stratum integrand is $g(x_{1})=\exp(-x_{1}^{2})\cos(x_{1})$. The integrand $f(\bm{z})$ for $\bm{z}\in[0,1]^{2}$ has a first component $z_{1}$ that selects a stratum $\ell$ as described in Section 2.2. Then $x_{1}=\theta_{\ell}+\Phi^{-1}(z_{2})$ where $\Phi(\cdot)$ is the cumulative distribution function of the $\mathcal{N}(0,1)$ distribution. We chose this function because it is regular enough that we expect to see an RQMC variance of $O(n^{-3})$.
We compare the variance of these unbiased estimators:
  1. 1)
    A plain Monte-Carlo estimator: $\hat{\mu}_{\mathrm{MC}}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathbf{U}(0,1)^{2}$.
  2. 2)
    A plain RQMC estimator: $\hat{\mu}_{\mathrm{RQMC}}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are the first $n$ points of a scrambled Sobol’ sequence in $[0,1]^{2}$.
  3. 3)
    An importance adjusted estimator: $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})\omega(\ell(i))$ with RQMC points where the $n_{\ell}$ come from rounding equation (14), using $\rho=2$, because that is the rate from the K-H bound for $n_{\ell}$ not equal to powers of two.
  4. 4)
    A reweighted RQMC estimator: $\hat{\mu}_{\mathrm{RQMC}(2)}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})\omega(\ell(i))$, where the $n_{\ell}$ are powers of $2$ chosen by Algorithm 1 with $\rho=3$, the variance rate for such $n_{\ell}$.
  5. 5)
    A per stratum estimator: $\hat{\mu}_{\mathrm{RQMC}(L)}=\sum_{\ell=1}^{L}\alpha_{\ell}\hat{\mu}_{\ell}$ where $\hat{\mu}_{\ell}$ are independently sampled RQMC estimates, with the same $n_{\ell}$ as in $\hat{\mu}_{\mathrm{RQMC}(2)}$.
For each estimator we compute $R=500$ replicates at sample sizes $n=2^{m}$ for $3\leqslant m\leqslant 12$ and then graph the sample variances.
The results are shown in Figure 2. The smaller sample sizes are comparable to the number of strata and so when we look for an understanding of convergence we emphasize the larger values of $n$. All of the methods are unbiased even though some of them have some empty strata at small $n$.
The variance for Monte Carlo sampling follows the expected $O(1/n)$ rate. Simply replacing the MC points by RQMC points improves the rate to roughly $O(1/n^{2})$. Adjusting the sampling fractions from $\alpha_{\ell}$ to $\beta_{\ell}$ using $\rho=2$ retains the RQMC rate and provides lower variance at larger samples sizes. Using $\rho=3$ to choose sample sizes $n_{\ell}$ we compared two methods. The estimate $\hat{\mu}_{\mathrm{RQMC}(L)}$ uses $L$ independent RQMC samples, with $n_{\ell}$ points in stratum $\ell$. The other uses the combined RQMC estimate $\hat{\mu}_{\mathrm{RQMC}(2)}$. These latter two both give an apparent $O(n^{-3})$ rate for the variance. The combined estimate has a better constant in this example. See our discussion of between strata correlations in Section 8.
Figure 2: Variance of the 5 estimators for the estimation of ${\mathbb{E}}(f(\bm{z}))$ versus the sample size $n$, for the toy integrand. Four of the curves are nearly parallel to reference lines of a given slope.
We also compared designs for different values of $\rho$ when taking $n_{\ell}$ to be powers of $2$. All of the MSEs were at the $n^{-3}$ rate. The best constant was from using $\rho=3$ and the worst was from $\rho=\infty$ which makes the $n_{\ell}$ equal. The variances for $\rho=1$ and $\rho=2$ were the same because those used the same values of $n_{\ell}=2^{m_{\ell}}$.
We noticed an unplanned anomaly in this example because we have $L=8$. Then using $\rho=\infty$, we get $n_{\ell}=n/8$, which is always a power of $2$ whether we use Algorithm 1 or just equation (14). As a result, using $\rho=\infty$ always gave $O(n^{-3})$ convergence where the best we could have had for $L=7$ or $9$ without forcing $n_{\ell}$ to be powers of two, would have been $O(n^{-2})$.
7 Saint-Venant flood model
We are motivated by engineering problems with substantial uncertainty. Great uncertainty often arises from weather variables or customer behavior. Our example is the Saint-Venant flood model from Limbourg and De Rocquigny (2010). A more precise flood model would be based on Navier-Stokes equations that have to be solved numerically. The Saint-Venant model is a simplification that we think is well suited for illustrative purposes.
The quantity of interest is the depth $H$ in meters of river water at a downstream location. The model has
$H=\Bigl{(}\frac{Q}{K_{s}B\sqrt{\alpha}}\Bigr{)}^{3/5},$ (20)
where $Q$ is the flow rate (meters3/second), $B$ is the width of the river (meters), $K_{s}$ is the Strickler coefficient (meters1/3/second) quantifying the roughness of the river bed and $\alpha$ is the (dimensionless) slope of the river bed. It is given by $\alpha=(Z_{m}-Z_{v})/L$ where $Z_{m}$ and $Z_{v}$ are the upstream and downstream heights of the river in meters.
Variable Nominal Adverse
$Q$ Fréchet (shape $6$, scale $1300$) Fréchet (shape $6$, scale $3900$)
$K_{s}$ Gamma (shape $90$, scale $1/3$) Gamma (shape $15$, scale $1$)
$Z_{v}$ Uniform ($49$,$51$) none
$Z_{m}$ Uniform ($54$,$56$) none
Table 2: Here are nominal distributions for four variables in the Saint-Venant model. Two of them also have adverse distributions that bring worse flooding.
The inputs to $H$ are random and prior researchers sampled each input independently from suitable distributions. Those prior contexts did not have mixtures. We have modified those distributions to include mixtures of nominal and adverse distributions. Our strata involve different subsets of the variables being adversely sampled.
This model is usually given with a symmetric triangular density for $L$ over the interval $[4990,5010]$. It is usually seen that uncertainty in $L$ makes very little difference. See for example Iooss and Lemaître (2015). Therefore we fix $L=5000$. Similarly, $B$ is commonly given a symmetric triangular distribution over $[295,305]$ and we fix it at $B=300$. We consider four distributions for the inputs $Q$, $K_{s}$, $Z_{v}$ and $Z_{m}$. With probability $0.95$ all four random variables are independently distributed with nominal distributions given in Table 2. With probability $0.02$, the distribution of $Q$ takes the adverse distribution in Table 2 while the other variables are nominal. With probability $0.02$, the distribution of $K_{s}$ is adverse and the others are nominal. With probability $0.01$ both $Q$ and $K_{s}$ are adverse.
The distribution of $K_{s}$ is customarily given as $\mathcal{N}(30,7.5^{2})$ and then truncated to a range such as $(0,\infty)$ in Limbourg and De Rocquigny (2010) or $(15,\infty)$ in Iooss (2011). We have chosen a nominal Gamma distribution, so we don’t need any truncation. That Gamma distribution has mean $30$ but a lower variance than $7.5^{2}$. The adverse distribution for the Gamma distribution takes lower values which bring more flooding as $K_{s}$ appears in the denominator of $H$. The distribution of $Q$ is customarily given as Gumbel with mode $1013$ and scale $558$. This was found by a maximum likelihood fit, reported in Limbourg and De Rocquigny (2010). The Gumbel distribution has an extreme value justification, but it also gives negative values which are unphysical. We have replaced it by a Fréchet distribution which is also an extreme value distribution. It has much heavier tails than the Gumbel which one might need to characterize extreme risks. Our nominal Fréchet distribution has a similar mean to the Gumbel distribution above. The adverse distribution triples the nominal one.
Figure 3: We plot variance versus sample size for $\hat{\mu}_{\mathrm{MC}}$, $\hat{\mu}_{\mathrm{RQMC}}$, $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$, $\hat{\mu}_{\mathrm{RQMC}(2)}$ and $\hat{\mu}_{\mathrm{RQMC}(L)}$ for the Saint-Venant flood depth (20). The sample size allocations for $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$, $\hat{\mu}_{\mathrm{RQMC}(2)}$, and $\hat{\mu}_{\mathrm{RQMC}(L)}$ used $\rho=2$.
The results are shown in Figure 3. This integrand is unbounded and hence not one for which the $O(n^{-3+\epsilon})$ results apply, and so we used $\rho=2$. We see a large gain from simply replacing MC points by RQMC points as in $\hat{\mu}_{\mathrm{RQMC}}$. The RQMC methods that oversampled the strata with smaller $\alpha_{\ell}$ had lower variance than the plain RQMC. The best of those was $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$ which used equation (14) without requiring $n_{\ell}$ to be powers of two. For $\rho=2$, we don’t expect a better rate from taking $n_{\ell}=2^{m_{\ell}}$. Sampling the $L$ strata independently had roughly the same variance as using the first component of $\bm{z}_{i}$ to select a stratum, where we had expected an advantage for the second method.
In Figure 4 we consider integer allocations with different values of $\rho$ using equation 14 to select $n_{\ell}$. For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope $-1.8$. The curve for $\rho=2$ which is also in Figure 3 is best at larger $n$. The case of $\rho=1$ has proportional allocation and may have an empty stratum for $n\leqslant 128$ (as $128*0.01<2$).
Figure 4: Variance of $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$ versus the sample size $n$ for $\rho\in\{1,2,3,+\infty\}$. The reference line has slope $-1.8$.
8 Discussion
We saw RQMC improving the sampling in a mixture model setting where there is one categorical variable. We found that taking account of the better than Monte Carlo convergence rates motivated changes where a stratum with mixture probability $\alpha_{\ell}$ would be better sampled with probability proportional to $\alpha_{\ell}^{\theta}$ for some $\theta<1$. Choosing the wrong power would adversely affect the constant in the asymptotic variance but not the rate.
A good exponent $\theta$ can be derived from an assumed convergence rate, and such a convergence rate can be estimated adaptively from pilot runs. We don’t define a formal adaptive algorithm. We are not sure of a way to do that which would be better than ad hoc experimentation.
We anticipate negative correlations among the $\hat{\mu}_{\ell}$ though we can also construct counter-examples. We expect that our within-stratum integrands $h_{\ell}(\bm{u})$ will tend to be similar from stratum to stratum in typical use cases. As a tiny example, suppose that $\alpha_{1}=\alpha_{2}=1/2$ and that $h_{1}(\bm{u})=h_{2}(\bm{u})=u_{1}$. The first $n/2$ values of $u_{1}$ (used in $h_{1}$) combine with the second $n/2$ values of $u_{1}$ (used in $h_{2}$) to give better uniformity than we get from either half on its own. We would get variance $1/(12n^{3})$ instead of $2/(12(n/2)^{3})=4/(3n^{3})$, which is sixteen times larger. A similar advantage would be expected when every $h_{\ell}$ is monotone in each $u_{j}$ either increasing in $u_{j}$ for all $\ell$ or decreasing in $u_{j}$ for all $\ell$. For example, increasing a physical variable might be harmful or helpful consistently across strata. We can of course induce positive correlations in the $\hat{\mu}_{\ell}$ by taking $h_{1}(\bm{u})=u_{1}$ and $h_{2}(\bm{u})=-u_{1}$. If one expects positive correlations then one might instead use independently sampled RQMC points in different strata. Independent sampling performed worse than conjoined sampling in the toy problem but the two were roughly equal in the Saint-Venant problem. We speculate that the negative correlation advantage, if it is present, is greatest when several $\beta_{\ell}$ are large.
We also tried a complex importance sampling problem from Hesterberg (1995) modeling reliability of oil inventory for an energy utility. The integrand was of the mixture importance sampling form (1). We found only a moderate improvement with RQMC giving roughly $O(n^{-1.15})$ variance. The example comes from Hesterberg (1995) and was also studied by Li et al. (2013). Here we only mention it to show that RQMC with mixture sampling will not always give an improvement that is practically meaningful. The relevant integrands were discontinuous and we speculate that they may have had important high order interactions.
We have considered one categorical variable. Some problems may well have two or more categorical variables. Suppose that there are $K$ categorical variables where the $k$’th one has $L_{k}$ levels. Then we have the choice of whether to use $K$ different components of RQMC to make the allocations, or to use one component with $L=\prod_{k=1}^{K}L_{k}$ levels in it. Optimizing the choice between these two approaches is outside the scope of this article.
Acknowledgments
We thank Neil Sloane for pointing out the connection to sequence A002572. Thanks to Tim Hesterberg and Wentao Li for help with the oil inventory problem. This work was supported by the National Science Foundation under grant DMS-2152780.
Appendix A Monotonocity of $I_{0}$
Here we prove Proposition 4.
Proof.
For fixed $\rho$ we may write
$I_{0}(\gamma\!\mid\!\rho)\propto\tilde{I}_{0}(\gamma\!\mid\!\rho)=\sum_{\ell=1}^{L}\alpha_{\ell}^{2-2\rho/(\gamma+1)}\Biggl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+1)}\Biggr{)}^{\rho},$
so it is enough to show that (17) holds with $I_{0}$ replaced by $\tilde{I}_{0}$. Now write $2/(\gamma+1)=2/(\rho+1)+x$, so we can use $x$ to characterize how $\gamma$ differs from $\rho$ with $x=0$ corresponding to $\gamma=\rho$. Then we can write $\tilde{I}_{0}(\gamma\!\mid\!\rho)=h(x)$ where
$h(x)$ $=\biggl{(}\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2-2\rho/(\rho+1)-\rho x}\biggr{)}\biggl{(}\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr{)}^{\rho}$
$=\biggl{(}\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2/(\rho+1)-\rho x}\biggr{)}\biggl{(}\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr{)}^{\rho}.$
It is enough to show that $h(x)$ is non-decreasing as $x$ either increases from zero or decreases from zero. We will do this by showing that $xh^{\prime}(x)\geqslant 0$.
For integers $r\geqslant 0$ define
$S_{+}(r)$ $=S_{+}(r,x)=\sum_{\ell=1}^{L}\alpha_{\ell}^{2/(\rho+1)+x}\log(\alpha_{\ell})^{r}\quad\text{and}$
$S_{-}(r)$ $=S_{-}(r,x)=\sum_{\ell=1}^{L}\alpha_{\ell}^{2/(\rho+1)-\rho x}\log(\alpha_{\ell})^{r}.$
Their derivatives with respect to $x$ are $S_{+}(r+1)$ and $-\rho S_{-}(r+1)$, respectively. Also, $S_{+}(r,0)=S_{-}(r,0)$ which we write as $S(r)$.
Now we write
$h(x)=S_{-}(0)S_{+}(0)^{\rho}$
so
$h^{\prime}(x)=-\rho S_{-}(1)S_{+}(0)^{\rho}+\rho S_{-}(0)S_{+}(0)^{\rho-1}S_{+}(1).$
Then
$\frac{h^{\prime}(x)}{\rho S_{+}(0)^{\rho-1}}$ $=S_{+}(1)S_{-}(0)-S_{+}(0)S_{-}(1)$
$=\sum_{i=1}^{L}\sum_{j=1}^{L}\alpha_{i}^{2/(\rho+1)-\rho x}\alpha_{j}^{2/(\rho+1)+x}\log(\alpha_{j})-\alpha_{i}^{2/(\rho+1)-\rho x}\alpha_{j}^{2/(\rho+1)+x}\log(\alpha_{i})$
$=\sum_{i=1}^{L-1}\sum_{j=i+1}^{L}\log\Bigl{(}\frac{\alpha_{j}}{\alpha_{i}}\Bigr{)}c_{ij}$
where
$c_{ij}$ $=\alpha_{i}^{2/(\rho+1)-\rho x}\alpha_{j}^{2/(\rho+1)+x}-\alpha_{i}^{2/(\rho+1)+x}\alpha_{j}^{2/(\rho+1)-\rho x}$
$=\alpha_{i}^{2/(\rho+1)}\alpha_{j}^{2/(\rho+1)}\Bigl{(}\Bigl{(}\frac{\alpha_{j}}{\alpha_{i}^{\rho}}\Bigr{)}^{x}-\Bigl{(}\frac{\alpha_{i}}{\alpha_{j}^{\rho}}\Bigr{)}^{x}\Bigr{)}.$
If $\alpha_{j}\geqslant\alpha_{i}$, then $\alpha_{j}/\alpha_{i}^{\rho}\geqslant\alpha_{i}/\alpha_{j}^{\rho}$ and so $xc_{ij}\geqslant 0$. Because $\log(\alpha_{j}/\alpha_{i})\geqslant 0$, we conclude that $x\log(\alpha_{j}/\alpha_{i})c_{ij}\geqslant 0$. If instead $\alpha_{j}\leqslant\alpha_{i}$, then a similar argument shows that $xc_{ij}\leqslant 0$ and once again $x\log(\alpha_{j}/\alpha_{i})c_{ij}\geqslant 0$. Hence
$xh^{\prime}(x)=\rho S_{+}(0)^{\rho-1}\sum_{i=1}^{L}\sum_{j=1}^{L}x\log(\alpha_{j}/\alpha_{i})c_{ij}\geqslant 0$
as required. ∎

### v2

@@
 We find that when integration error decreases at an RQMC rate that it is then beneficial to oversample the smallest mixture components instead of using a proportional allocation.
+The optimal allocations depend on the possibly unknown convergence rates.
+Designing the sample with an incorrect assumption on the rate still attains that convergence rate, with an inferior implied constant.
+The penalty for using a pessimistic rate is typically higher than for using an optimistic one.
 We also find that for the most accurate RQMC sampling methods, it is advantageous to arrange that our $n=2^{m}$ randomized Sobol’ points split into subsample sizes that are also powers of $2$.
@@
 The main motivation for this problem is mixture sampling where the variables of interest come from $L$ different distributions representing $L$ different scenarios.
-That is $\bm{x}\sim P_{\ell}$ with probability $\alpha_{\ell}>0$ where $\sum_{\ell=1}^{L}\alpha_{\ell}=1$.
+That is, $\bm{x}\sim P_{\ell}$ with probability $\alpha_{\ell}>0$ where $\sum_{\ell=1}^{L}\alpha_{\ell}=1$.
 The distributions $P_{\ell}$ are commonly called mixture components.
 We call them ‘strata’ to avoid confusion with vector components.
+We want to estimate $\mu={\mathbb{E}}(g(\bm{x}))=\sum_{\ell=1}^{L}\alpha_{\ell}{\mathbb{E}}(g(\bm{x})\!\mid\!\bm{x}\sim P_{\ell})$.
+Our approach will also allow us to use distinct functions $g_{\ell}$ within each stratum.
 A straightforward approach to this problem is to simply replace a standard Monte Carlo (MC) workflow with RQMC sampling.
@@
 That approach misses a potential variance reduction from RQMC sampling when the within-stratum estimates are negatively correlated.
-Such negative correlation need not happen; we discuss them briefly at the end of the paper.
-For some problems the integrand is a simple function within each stratum.
+Such negative correlation need not happen; we give a reason to expect that at the end of the paper.
+For some problems, the integrand is a simple function within each stratum.
 A more interesting expression arises in mixture importance sampling.
@@
 Such a rate can only be attained over a sequence of special sample sizes that grow at least exponentially (Sobol’, 1998; Owen, 2016).
-This rate can be attained at powers of $2$ by scrambled Sobol’ points.
+This rate can be attained by scrambled Sobol’ points with sample size equal to a power of $2$.
 Even an integrand with a QMC-friendly axis-parallel discrepancy is not generally integrated with an error of this rate.
 By using an importance sampling (Owen, 2013, Chapter 9) adjustment, we can arrange to obtain $n_{\ell}=2^{m_{\ell}}$ points from $P_{\ell}$ while also having $n=\sum_{\ell=1}^{L}n_{\ell}$ be a power of $2$.
-This can then give the $O(n_{\ell}^{-3+\epsilon})$ rate within stratum $\ell$ and hence overall.
+This can then give the $O(n_{\ell}^{-3+\epsilon})$ rate within every stratum and hence overall.
 This paper is organized as follows.
@@
 Section 3 describes how to allocate samples to strata.
-In RQMC sampling, it is advantageous to use more nearly equal allocations than those given by the stratum sizes.
+We see that in RQMC sampling, it is advantageous to use more nearly equal allocations than those given by the stratum sizes.
 Section 4 shows how to partition a sample size $n=2^{m}$ into $L$ within-stratum sample sizes $n_{\ell}=2^{m_{\ell}}$.
@@
 Section 6 shows a simple toy example with $L=8$ strata where RQMC with $n_{\ell}$ taken to be powers of two does best.
-Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg and De Rocquigny (2010) where non-nominal sample sizes do best.
+Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg and De Rocquigny (2010) where non-nominal sampling proportions do best.
 Section 8 has some discussion and conclusions.
@@
 They construct a continuous mapping over $t\in[0,1]$ with $\mathcal{N}(0,I)$ at $t=0$ reaching $P$ at $t=1$ and solve an ordinary differential equation along that path.
-For $\ell<L$ they take the first $n_{\ell}=\lfloor n\alpha_{\ell}\rfloor$ points of a quasi-Monte Carlo sequence and use those to sample from component $\ell$.
+For $\ell<L$, they take the first $n_{\ell}=\lfloor n\alpha_{\ell}\rfloor$ points of a quasi-Monte Carlo sequence and use those to sample from component $\ell$.
 Then $n_{L}=n-\sum_{\ell=1}^{L-1}n_{\ell}$.
@@
 We will use RQMC points $\bm{z}\in[0,1]^{s+1}$.
-We partition $\bm{z}$ into its first component $v$ that we use to select a stratum $\ell$ and the subsequent $s$ components $\bm{u}$ that we use to sample $\bm{x}$ from $P_{\ell}$.
+We partition $\bm{z}$ into its first component $v$ that we use to select a stratum $\ell$ and the subsequent $s$ components $\bm{u}$ that we use to sample $\bm{x}\sim P_{\ell}$.
 That is, $\bm{z}=(v,u_{1},u_{2},\dots,u_{s})$.
@@
 is a within-stratum sample mean.
-For the RQMC method we consider we show below that $n_{\ell}\geqslant 1$ whenever $\beta_{\ell}>2/n$.
-We do not need $\Pr(n_{\ell}>0)=1$ for unbiasedness of $\hat{\mu}$, but without it the stratum sample means are not always well defined.
+For the RQMC method we consider, we show below that $n_{\ell}\geqslant 1$ whenever $\beta_{\ell}>2/n$.
+We do not need $\Pr(n_{\ell}>0)=1$ for unbiasedness of $\hat{\mu}$, but without it the stratum sample means $\hat{\mu}_{\ell}$ are not always well defined.
 At this point we can introduce a small generalization.
-Since we already have a dependence on $\ell$ in $h_{\ell}$ we could also replace $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))$ by $h_{\ell}(\bm{u})=g_{\ell}(\phi_{\ell}(\bm{u}))$ to allow the function being applied to $\bm{x}$ to vary with $\ell$.
+Since we already have a dependence on $\ell$ in $h_{\ell}$ we could also replace $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))$ by $h_{\ell}(\bm{u})=g_{\ell}(\phi_{\ell}(\bm{u}))=g_{\ell}(\bm{x})$ to allow the function being applied to $\bm{x}$ to vary with $\ell$.
 2.2 Sampling the strata
@@
 By this we mean that every interval $I_{k}=[(k-1)/n,k/n)$ contains precisely one of $v_{1},\dots,v_{n}$ for $k=1,\dots,n$.
-The next proposition shows that if every $\beta_{\ell}>2/n$ then every $n_{\ell}>0$ when $v_{i}$ are stratified.
+The next proposition shows that if every $\beta_{\ell}>2/n$, then every $n_{\ell}>0$ when $v_{i}$ are stratified.
 Proposition 1.
@@
 Write $A=r/n+a$ and $B=s/n+b$ for integers $0\leqslant r\leqslant s<n$ and $a,b\in[0,1/n)$.
-Then $\beta=t/n+c$ for $t=r-s$ and $-1/n<c<1/n$.
+Then $\beta=t/n+c$ for $t=s-r$ and $-1/n<c<1/n$.
 If $r=s$, then $0\leqslant n_{*}\leqslant 1$ and then (6) holds trivially.
@@
 for integers $k_{j}\geqslant 0$ and $0\leqslant c_{j}<2^{k_{j}}$. We will use $|\bm{k}|=\sum_{j=1}^{d}k_{j}$, so $H(\bm{c},\bm{k})$ has volume $2^{-|\bm{k}|}$.
-For integers $m\geqslant t\geqslant 0$ the $n=2^{m}$ points $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1)^{d}$ are a $(t,m,d)$-net in base $b$ if $H(\bm{c},\bm{k})$ contains precisely $nb^{-|\bm{k}|}=b^{m-|\bm{k}|}$ of them, whenever $m-|\bm{k}|\geqslant t$.
+For integers $m\geqslant t\geqslant 0$ the $n=b^{m}$ points $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1)^{d}$ are a $(t,m,d)$-net in base $b$ if $H(\bm{c},\bm{k})$ contains precisely $nb^{-|\bm{k}|}=b^{m-|\bm{k}|}$ of them, whenever $m-|\bm{k}|\geqslant t$.
 Smaller $t$ are better.
@@
 For a discussion of that point, see Owen and Pan (2022).
+While convergence $O(n^{-\rho})$ for $\rho\in\{1,2,3\}$ are widely studied, these come from error bounds and the actual convergence rate need not be near an integer.
+Liu (2025) shows examples where certain singular behavior of the integrand can give any rate between $1$ and $2$.
+In applications, the precise rate of convergence might not be known before computation begins.
 2.5 Discrepancy within each stratum
@@
 Then the points in a given stratum still have small discrepancy but they are not generally scrambled digital nets.
-In that case the rate in stratum $\ell$ is $O(n_{\ell}^{-2+\epsilon})$ for within-stratum integrands that are in BVHK.
+In that case the variance rate in stratum $\ell$ is $O(n_{\ell}^{-2+\epsilon})$ for within-stratum integrands that are in BVHK.
 Proposition 2.
@@
 If $m-\kappa_{\ell}-|\bm{k}|\geqslant t$, then by the digital net property of $\bm{z}_{1},\dots,\bm{z}_{n}$, this number is $2^{m-\kappa_{\ell}-|\bm{k}|}$.
-That is, if $m_{\ell}\geqslant|\bm{k}|+t$, then the number of such points is $2^{m_{\ell}-|\bm{k}|}$
+That is, if $m_{\ell}\geqslant|\bm{k}|+t$, then the number of such points is $2^{m_{\ell}-|\bm{k}|}$.
+∎
 To motivate the constraint that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$, consider what would happen if we took $\beta_{1}=1/8$, $\beta_{2}=1/2$, $\beta_{3}=1/4$ and $\beta_{4}=1/8$.
@@
 and so the star discrepancy of $\bm{u}_{i}$ for $i\in S_{\ell}$ is $O(n_{\ell}^{-1}\log(n_{\ell})^{s})$. ∎
-If $\bm{z}_{1},\dots,\bm{z}_{n}$ are Halton points (scrambled or otherwise) with $n=2^{m}$ then $\bm{z}_{i1}$ for $1\leqslant i\leqslant n$ are stratified.
 3 Choosing $\beta_{\ell}$ for given $\alpha_{\ell}$
@@
 We will show that choosing $\beta_{\ell}$ based on the wrong value of $\rho$ can adversely affect the implied constant in the asymptotic MSE while leaving the convergence rate unchanged.
-Values of $\rho$ equal to $1$, $2$ and $3$ correspond to upper bounds on the covergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases.
+Values of $\rho$ equal to $1$, $2$ and $3$ correspond to upper bounds on the convergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases.
 It often happens that the empirical MSE of an RQMC rule follows an apparent rate of $n^{-\rho}$ for a value of $\rho$ that is not close to an integer.
-That can happen when the practically usable values of $n$ are smaller than the sample size at which the asymptotic result holds.
 Unbounded integrands are not in BVHK, and yet they will have an MSE that is $o(n^{-1})$.
@@
 Section 4 shows how to obtain $n_{\ell}$ equal to powers of two.
-With $n_{\ell}=n\beta_{\ell}$ and $\omega_{\ell}=\alpha_{\ell}/\beta_{\ell}$, we can rewrite (5) as
+Either way, we can assume that $n_{\ell}=n\beta_{\ell}$. Then with $\omega_{\ell}=\alpha_{\ell}/\beta_{\ell}$, we can rewrite (5) as
 $\hat{\mu}=\sum_{\ell=1}^{L}\omega_{\ell}\frac{n_{\ell}}{n}\hat{\mu}_{\ell}=\sum_{\ell=1}^{L}\alpha_{\ell}\hat{\mu}_{\ell}.$
-The importance sampling ratio $\omega_{\ell}$ has canceled from the expression for $\hat{\mu}$, leaving a simply weighted some of within-stratum means.
+The importance sampling ratio $\omega_{\ell}$ has canceled from the expression for $\hat{\mu}$, leaving a simply weighted sum of within-stratum means.
 3.1 Optimal sample sizes
@@
 RQMC points have a space filling property.
-We expect that the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$.
+By construction, the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$.
 We might then anticipate primarily non-positive correlations among the $\hat{\mu}_{\ell}$.
@@
 $n_{\ell}\propto\alpha_{\ell}^{2/(\rho+1)}.$ (14)
-Higher convergence rates $\rho$ tend to equalize the $n_{\ell}$.
-We can think of $n_{\ell}=n/L$ as arising in an $\rho\to\infty$ limit.
+In both criteria (12) and (14), higher convergence rates $\rho$ tend to equalize the $n_{\ell}$.
+We can think of $n_{\ell}=n/L$ as arising in the limit as $\rho\to\infty$.
 We will also see equal $n_{\ell}$ in a minimax argument in Section 5.
-We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for ansatzs 00 and $1$, and three different hypotheses on $\rho$.
+We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for ansatzes 00 and $1$, and three different hypotheses on $\rho$.
 They lead to the four different rates given in Table 1.
@@
 The other choices all allocate more than a proportional number of samples to the strata with smaller $\alpha_{\ell}$.
-This is even true for the plain MC rate ($\rho=1$) under the conservative ansatz 1.
-When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then equations (12) and (14) don’t distinguish betweem different values of $\rho$.
+This is also true for the plain MC rate ($\rho=1$) under the conservative ansatz 1.
+When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then equations (12) and (14) do not distinguish between different values of $\rho$.
 After raising $\alpha_{\ell}$ to any power they are still equal and then a near equal integer allocation is appropriate.
@@
 In that case, the proportionality for $C_{0}$ and $\rho=1$ matches the Neyman allocation from survey sampling, in Neyman (1934).
-When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$ then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}(\tau_{\ell}/c_{\ell})^{1/2}$ in Table 1.
+When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$, then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}/c_{\ell}$ in Table 1.
 $\rho:$ 1 2 3
@@
 These rates apply when the variance within stratum $\ell$ decreases proportionally to $n_{\ell}^{-\rho}$.
-If $\rho$ varies from stratum, taking the value $\rho_{\ell}$ in stratum $\ell$, then the Lagrange multiplier argument leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho_{\ell}+1)}$.
+If $\rho$ varies from stratum to stratum, taking the value $\rho_{\ell}$ in stratum $\ell$, then the Lagrange multiplier argument leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho_{\ell}+1)}$.
 Strata with higher $\rho_{\ell}$ get larger $n_{\ell}$ and $\beta_{\ell}=n_{\ell}/n$ does not depend on $n$.
@@
 We measure inefficiency by the ratio of the MSE we get if we design with a value $\gamma$ when we should have used $\rho$, so smaller values are better, and $1$ is best possible.
-In ansatz 0 we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
+In ansatz 0, we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
 $I_{0}(\gamma\!\mid\!\rho)$ $=\frac{\sum_{\ell=1}^{L}\alpha^{2}_{\ell}\Bigl{(}n\alpha_{\ell}^{2/(\gamma+1)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+1)}\Bigr{)}^{-\rho}}{\sum_{\ell=1}^{L}\alpha^{2}_{\ell}\Bigl{(}n\alpha_{\ell}^{2/(\rho+1)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)}\Bigr{)}^{-\rho}}$
@@
 When we know that $\rho\leqslant 2$ we might only let $\rho$ and $\gamma$ belong to $[1,2]$.
-The minimax approach of equations (16) gives us ways to choose $\gamma$ for a given list of values $\alpha_{1},\dots,\alpha_{L}$.
-It often happens that the worst value of $\rho$ for given $\gamma$ is one of the allowed extremes, but this does not always happen.
+The minimax approach of equation (16) gives us ways to choose $\gamma$ for a given list of values $\alpha_{1},\dots,\alpha_{L}$.
+It often happens that the worst value of $\rho$ for given $\gamma$ is at one of the allowed extremes, but this does not always happen.
 We have seen such exceptions when one of the $\alpha_{\ell}$ is much larger than the others.
@@
 After rearrangement this becomes a constraint that $\gamma\geqslant\rho-1$, whatever the unknown $\rho$ might be.
-When $\rho\doteq 3$ is a possibility then we should take $\gamma\geqslant 2$.
+When $\rho\doteq 3$ is a possibility, then we should take $\gamma\geqslant 2$.
 If instead we use $\gamma=1$ in a $\rho=3$ setting then the first factor in $I_{0}$ is greater than $\alpha_{L}^{-1}$.
@@
 4 Making each $n_{\ell}$ a power of two
-Here we consider how to allocate sample sizes $n_{\ell}=2^{m_{\ell}}$ that sum to given a total sample size $n=2^{m}$.
+Here we consider how to allocate sample sizes $n_{\ell}=2^{m_{\ell}}$ that sum to a given total sample size $n=2^{m}$.
 We assume throughout that $n\geqslant L$ and that we want every $n_{\ell}\geqslant 1$.
@@
 The results with MSE $O(n^{-3+\epsilon})$ for scrambled Sobol’ points require such sample sizes.
-The potential disadvantage of having $n_{\ell}$ be a power of $2$ is that we have many fewer sampling ratios $\alpha_{\ell}/\beta_{\ell}$ at our disposal.
+A potential non-asymptotic disadvantage of having $n_{\ell}$ be a power of $2$ is that we have many fewer sampling ratios $\alpha_{\ell}/\beta_{\ell}$ at our disposal.
 For example, with $L=2$, we would have to use $n_{1}=n_{2}=n/2$ and this could be a bad fit for $\alpha_{1}=0.9999$ and $\alpha_{2}=0.0001$.
@@
 For $L=2$ the only solution is $(1/2,1/2)$ and for $L=3$ the only solution is $(1/2,1/4,1/4)$.
-For $L=4$, there are two possibilities: $(1/2,1/4,1/8,1/8)$ and $(1/4,1/4,1/4,1/4)$ and for $5\leqslant L\leqslant 10$ the numbers of possibilities are: 3, 5, 9, 16, and 28.
+For $L=4$, there are two possibilities: $(1/2,1/4,1/8,1/8)$ and $(1/4,1/4,1/4,1/4)$ and for $5\leqslant L\leqslant 10$ the numbers of possibilities are: 3, 5, 9, 16, 28 and 50.
 The tables in Figure 1 show the partitions for $4\leqslant L\leqslant 8$.
@@
 This is a somewhat faster growth rate than the Fibonacci numbers have: their counterpart to $\theta$ is $(1+\sqrt{5})/2\doteq 1.618$.
-That asymptote is not a good description for small $L$.
-For $L=10$, it is over 4000, while the exact value is 28.
-To choose $\beta_{\ell}$, or equivalently $n_{\ell}$, we uses a forward stepwise algorithm.
+For $L=10$, the approximation is close to 48.3.
+To choose $\beta_{\ell}$, or equivalently $n_{\ell}$, we use a forward stepwise algorithm.
 We initialize every $m_{\ell}=0$ so every $n_{\ell}=2^{m_{\ell}}$ is initially $1$.
@@
 Proof.
-First, if $n^{\prime}>0$ then there is an ‘eligible’ index $\ell$ with $n_{\ell}\leqslant n^{\prime}$, that can be doubled without causing $\sum_{\ell=1}^{L}n_{\ell}>n$.
+First, if $n^{\prime}>0$ then we can show that there is an ‘eligible’ index $\ell$ with $n_{\ell}\leqslant n^{\prime}$, that can be doubled without causing $\sum_{\ell=1}^{L}n_{\ell}>n$.
 At any stage in the algorithm, $n_{\ell^{*}}$ for $\ell^{*}=\operatorname*{argmin}_{1\leqslant\ell\leqslant L}n_{\ell}$ is a power of two.
-It divides $n$ and it divides every $n_{\ell}$, so it divides the nonzero value $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$, hence $n_{\ell^{*}}\leqslant n^{\prime}$.
+It divides $n$ and it divides every $n_{\ell}$, so it also divides the nonzero value $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$, hence $n_{\ell^{*}}\leqslant n^{\prime}$.
 Finally, every step with $n^{\prime}>0$ reduces $n^{\prime}$, and so $n^{\prime}$ must reach zero in a finite number of steps starting from $n^{\prime}=n-L$.
@@
 We consider the increase in variance that we get using sample sizes $\bm{n}$ compared to some better sample sizes $\tilde{\bm{n}}$ that we might otherwise have used.
-For ansatzes $a=0,1$ we define suboptimality ratios
-$R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}}{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}\tilde{n}_{\ell}^{-\rho/2}}\quad\text{and}\quad R_{0}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}n_{\ell}^{-\rho}}{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}\tilde{n}_{\ell}^{-\rho}}$
+For ansatzes 00 and $1$ we define suboptimality ratios
+$R_{0}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}n_{\ell}^{-\rho}}{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}\tilde{n}_{\ell}^{-\rho}}\quad\text{and}\quad R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}}{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}\tilde{n}_{\ell}^{-\rho/2}}$
 that compare different sample sizes taking account of the stratum variances $\tau_{\ell}$.
@@
 We chose this function because it is regular enough that we expect to see an RQMC variance of $O(n^{-3})$.
+We chose these values of $\alpha_{\ell}$ so that the forward algorithm would give different $n_{\ell}$ for $\rho=2$ than it does for $\rho=3$.
 We compare the variance of these unbiased estimators:
@@
   3. 3)
-    An importance adjusted estimator: $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})\omega(\ell(i))$ with RQMC points where the $n_{\ell}$ come from rounding equation (14), using $\rho=2$, because that is the rate from the K-H bound for $n_{\ell}$ not equal to powers of two.
+    An importance adjusted estimator: $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})\omega(\ell(i))$ with RQMC points where the $n_{\ell}$ come from rounding equation (14), using $\rho=2$, because that is the rate from the Koksma-Hlawka bound for $n_{\ell}$ not equal to powers of two.
   4. 4)
@@
 For each estimator we compute $R=500$ replicates at sample sizes $n=2^{m}$ for $3\leqslant m\leqslant 12$ and then graph the sample variances.
+We used nested uniform scrambling (Owen, 1995) to do the scrambling.
 The results are shown in Figure 2.
@@
 Adjusting the sampling fractions from $\alpha_{\ell}$ to $\beta_{\ell}$ using $\rho=2$ retains the RQMC rate and provides lower variance at larger samples sizes.
-Using $\rho=3$ to choose sample sizes $n_{\ell}$ we compared two methods.
-The estimate $\hat{\mu}_{\mathrm{RQMC}(L)}$ uses $L$ independent RQMC samples, with $n_{\ell}$ points in stratum $\ell$.
+Using $\rho=3$ to choose sample sizes $n_{\ell}$, we compared two methods.
+The first generates the estimate $\hat{\mu}_{\mathrm{RQMC}(L)}$ using $L$ independent RQMC samples, with $n_{\ell}$ points in stratum $\ell$.
 The other uses the combined RQMC estimate $\hat{\mu}_{\mathrm{RQMC}(2)}$.
@@
 where $Q$ is the flow rate (meters3/second), $B$ is the width of the river (meters), $K_{s}$ is the Strickler coefficient (meters1/3/second) quantifying the roughness of the river bed and $\alpha$ is the (dimensionless) slope of the river bed.
-It is given by $\alpha=(Z_{m}-Z_{v})/L$ where $Z_{m}$ and $Z_{v}$ are the upstream and downstream heights of the river in meters.
+It is given by $\alpha=(Z_{m}-Z_{v})/D$ where $Z_{m}$ and $Z_{v}$ are the upstream and downstream heights of the river in meters, and $D$ denotes the river’s length in meters.
 Variable Nominal Adverse
@@
 Our strata involve different subsets of the variables being adversely sampled.
-This model is usually given with a symmetric triangular density for $L$ over the interval $[4990,5010]$.
-It is usually seen that uncertainty in $L$ makes very little difference.
+This model is usually given with a symmetric triangular density for $D$ over the interval $[4990,5010]$.
+It is usually seen that uncertainty in $D$ makes very little difference.
 See for example Iooss and Lemaître (2015).
-Therefore we fix $L=5000$.
+Therefore we fix $D=5000$.
 Similarly, $B$ is commonly given a symmetric triangular distribution over $[295,305]$ and we fix it at $B=300$.
@@
 The distribution of $K_{s}$ is customarily given as $\mathcal{N}(30,7.5^{2})$ and then truncated to a range such as $(0,\infty)$ in Limbourg and De Rocquigny (2010) or $(15,\infty)$ in Iooss (2011).
-We have chosen a nominal Gamma distribution, so we don’t need any truncation.
+We have chosen a nominal Gamma distribution, so we do not need any truncation.
 That Gamma distribution has mean $30$ but a lower variance than $7.5^{2}$.
@@
 The Gumbel distribution has an extreme value justification, but it also gives negative values which are unphysical.
-We have replaced it by a Fréchet distribution which is also an extreme value distribution.
+We have replaced it by a Fréchet distribution which is always positive and is also an extreme value distribution.
 It has much heavier tails than the Gumbel which one might need to characterize extreme risks.
@@
 Sampling the $L$ strata independently had roughly the same variance as using the first component of $\bm{z}_{i}$ to select a stratum, where we had expected an advantage for the second method.
-In Figure 4 we consider integer allocations with different values of $\rho$ using equation 14 to select $n_{\ell}$.
-For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope $-1.8$.
+In Figure 4 we consider integer allocations with different values of $\rho$ using equation (14) to select $n_{\ell}$.
+For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope close to $-1.8$.
 The curve for $\rho=2$ which is also in Figure 3 is best at larger $n$.
@@
 The relevant integrands were discontinuous and we speculate that they may have had important high order interactions.
+We might also have a variable with a mixed distribution including a continuous component with a density function and $L$ atoms.
+For this setting we can make the continuous part of that variable’s distribution into an $L+1$’st mixture component.
+We will still need to sample within the range of that continuous component.
+For that we can add an $s+1$’st component to $\bm{u}$, which is then the $s+2$’nd component of $\bm{z}$.
 We have considered one categorical variable.

### v3

@@
 We study randomized quasi-Monte Carlo (RQMC) integration in a setting where one of the input variables is categorical, taking only some finite number $L\geqslant 2$ of different values.
-The main motivation for this problem is mixture sampling where the variables of interest come from $L$ different distributions representing $L$ different scenarios.
+The main motivation is mixture sampling where the variables of interest come from $L$ different distributions representing $L$ different scenarios.
 That is, $\bm{x}\sim P_{\ell}$ with probability $\alpha_{\ell}>0$ where $\sum_{\ell=1}^{L}\alpha_{\ell}=1$.
-The distributions $P_{\ell}$ are commonly called mixture components.
-We call them ‘strata’ to avoid confusion with vector components.
 We want to estimate $\mu={\mathbb{E}}(g(\bm{x}))=\sum_{\ell=1}^{L}\alpha_{\ell}{\mathbb{E}}(g(\bm{x})\!\mid\!\bm{x}\sim P_{\ell})$.
-Our approach will also allow us to use distinct functions $g_{\ell}$ within each stratum.
 A straightforward approach to this problem is to simply replace a standard Monte Carlo (MC) workflow with RQMC sampling.
-In such a workflow, one random variable selects stratum $\ell$ with probability $\alpha_{\ell}$.
-Then subsequent random variables are used to generate $\bm{x}\sim P_{\ell}$.
-When MC is replaced by RQMC, we find some advantages to sampling the $L$ strata more equally than their nominal proportions $\alpha_{\ell}$ would indicate.
+In that workflow, one random variable selects component $\ell$ with probability $\alpha_{\ell}$.
+Then $s$ more random variables generate $\bm{x}\sim P_{\ell}$.
+RQMC estimates over $s+1$ dimensions converge at a better rate than MC estimates do.
+Then the optimal sampling allocations become more nearly equal than proportional sampling, and we modify RQMC to take advantage of this phenomenon.
+We see reduced variance with those choices, and under regularity conditions, non-proportional sampling of the $P_{\ell}$ can even improve the convergence rate.
+We also show a minimaxity property for equal or nearly equal sampling fractions.
+The input points leading to $\bm{x}\sim P_{\ell}$ come from $L$ disjoint subsets or strata in $0,1]^{s+1}$.
+From here on we refer to the mixture components as strata to avoid confusion with vector components.
 An alternative to the standard MC workflow is to compute $L$ different independently sampled RQMC estimates, one per stratum, and combine them according to their mixture probabilities $\alpha_{\ell}$.
 That approach misses a potential variance reduction from RQMC sampling when the within-stratum estimates are negatively correlated.
-Such negative correlation need not happen; we give a reason to expect that at the end of the paper.
-For some problems, the integrand is a simple function within each stratum.
-A more interesting expression arises in mixture importance sampling.
-There we estimate $\mu=\int g(\bm{x})p(\bm{x})\mathrm{\,d}\bm{x}$ for a probability density function $p$ by
-$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\frac{g(\bm{x}_{i})p(\bm{x}_{i})}{\sum_{\ell=1}^{L}\alpha_{\ell}p_{\ell}(\bm{x}_{i})}$ (1)
-for $\bm{x}_{i}\sim\sum_{\ell=1}^{L}\alpha_{\ell}p_{\ell}$. By Jensen’s inequality, this has lower Monte Carlo variance than we would get from
-$\sum_{\ell=1}^{L}\frac{\alpha_{\ell}}{n_{\ell}}\sum_{i=1}^{n_{\ell}}\frac{g(\bm{x}_{i,\ell})p(\bm{x}_{i,\ell})}{p_{\ell}(\bm{x}_{i,\ell})}.$
-working separately within each of the $L$ strata.
+Such negative correlation need not happen; we give a reason to expect them at the end of the paper.
 RQMC does best under favorable regularity assumptions, such as continuity, on the integrand.
 The presence of the categorical variable makes the resulting integrand discontinuous.
-That discontinuity is however axis parallel and so it is a QMC-friendly discontinuity in the terminology of Wang and Tan (2013).
+That discontinuity is however axis-parallel and so it is a QMC-friendly discontinuity in the terminology of Wang & Tan (2013).
 In some very regular cases, RQMC can attain a mean squared error (MSE) of $O(n^{-3+\epsilon})$.
-Such a rate can only be attained over a sequence of special sample sizes that grow at least exponentially (Sobol’, 1998; Owen, 2016).
+Such a rate can only be attained over a sequence of special sample sizes that grow at least exponentially (Sobol’ 1998, Owen 2016).
 This rate can be attained by scrambled Sobol’ points with sample size equal to a power of $2$.
 Even an integrand with a QMC-friendly axis-parallel discrepancy is not generally integrated with an error of this rate.
-By using an importance sampling (Owen, 2013, Chapter 9) adjustment, we can arrange to obtain $n_{\ell}=2^{m_{\ell}}$ points from $P_{\ell}$ while also having $n=\sum_{\ell=1}^{L}n_{\ell}$ be a power of $2$.
+By using an importance sampling (Owen 2013, Chapter 9) adjustment, we can arrange to obtain $n_{\ell}=2^{m_{\ell}}$ points from $P_{\ell}$ while also having $n=\sum_{\ell=1}^{L}n_{\ell}$ be a power of $2$.
 This can then give the $O(n_{\ell}^{-3+\epsilon})$ rate within every stratum and hence overall.
@@
 Section 6 shows a simple toy example with $L=8$ strata where RQMC with $n_{\ell}$ taken to be powers of two does best.
-Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg and De Rocquigny (2010) where non-nominal sampling proportions do best.
+Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg & De Rocquigny (2010) where non-nominal sampling proportions do best.
+The code used to produce the results for both examples is publicly available at https://github.com/hoval58/qmc4categorical.
 Section 8 has some discussion and conclusions.
-The proof of one of our results is in Appendix A.
+There is an appendix for the proofs.
 We finish this introduction by describing some related work in the literature.
-Our discussion uses the notation of this paper to describe the findings of the other papers.
-The key differences between our work and the prior work we found is in how the problem has been formulated, and how the values are sampled.
+The key differences between our work and the prior work we found are in how the problem has been formulated, and how the values are sampled.
 We use the first component of our RQMC points to select the stratum to sample from.
-Keller and Grünschloß (2012) used that approach in parallel computation over $L$ processors.
+Keller & Grünschloß (2012) used that approach in parallel computation.
 Their first component was used to determine which processor $\ell\in\{1,2,\dots,L\}$ got the $i$’th job.
@@
 Cui et al.
-(2025) consider a target distribution on a bounded hyperrectangle $[\bm{a},\bm{b}]\subset^{d}$.
+(2025) consider a target distribution on a bounded hyperrectangle $\bm{a},\bm{b}]\subset\real^{d}$.
 The target probability density function $\pi$ is bounded.
@@
 They use the first $n_{j}$ points of the same QMC point set for each of the strata.
-Their integrands $f$ belong to an $L_{p}$ weighted space with a slightly different norm from the usual ones.
-Their equation (3) defining the norm also uses some sup-norms of absolute mixed partial derivatives.
-Then sampling on a class of digital nets, they are able to bound the integration error by an expression using the integrand’s norm.
-They have another bound when the density satisfies a Hölder condition.
 We use different RQMC points in each stratum.
@@
 Since our analysis is based on variances, we do not require a bounded domain or integrand.
-We also find an advantage, theoretically and empirically, to diverging from proportional allocation.
-Klebanov and Sullivan (2023) describe using transport maps from a $\mathcal{N}(0,I)$ reference distribution to sample a mixture like $P=\sum_{\ell=1}^{L}\alpha_{j}\mathcal{N}(\mu_{j},\Sigma_{j})$.
+As mentioned above, we use non-proportional allocation.
+Klebanov & Sullivan (2026) describe using transport maps from a $\mathcal{N}(0,I)$ reference distribution to sample a mixture like $P=\sum_{\ell=1}^{L}\alpha_{j}\mathcal{N}(\mu_{j},\Sigma_{j})$.
 They construct a continuous mapping over $t\in[0,1]$ with $\mathcal{N}(0,I)$ at $t=0$ reaching $P$ at $t=1$ and solve an ordinary differential equation along that path.
@@
 Then $n_{L}=n-\sum_{\ell=1}^{L-1}n_{\ell}$.
-This is different from our strategy of using the first RQMC component to select the stratum which we prefer because it avoids giving two within-stratum samplers the same points.
-They do not require $n\gg L$.
-Some of our expressions only require $n>2L$, but as we use asymptotic variances within strata, our results are most relevant for large $n/L$.
-They consider more general location-scale families than the Gaussian distribution and they can even include some $\alpha_{\ell}<0$ so long as $P$ never has a negative probability density function.
 2 Background
@@
 Then we define digital nets and discuss their randomizations.
-For a broader survey of RQMC see L’Ecuyer and Lemieux (2002).
-We use $\mathbb{N}$ for the set of positive integers. We use $\bm{1}\{A\}$ for the indicator function that equals $1$ when $A$ holds and is 00 otherwise.
+For a broader survey of RQMC see L’Ecuyer & Lemieux (2002).
+We use $\mathbb{N}$ for the set of positive integers. We use $\bm{1}\{A\}$ for the indicator function that equals $1$ when $A$ holds and is 0 otherwise.
 2.1 Integration variables
-Our goal is to estimate
-$\mu=\sum_{\ell=1}^{L}\alpha_{\ell}\int_{{}^{d}}g(\bm{x})p_{\ell}(\bm{x})\mathrm{\,d}\bm{x}$
+Our goal is to estimate this weighted sum of $d$-dimensional integrals
+$\mu=\sum_{\ell=1}^{L}\alpha_{\ell}\int_{\real{}^{d}}g(\bm{x})p_{\ell}(\bm{x})\mathrm{\,d}\bm{x}$
 where $p_{\ell}$ is the probability density function of $P_{\ell}$ and $g(\cdot)$ is the integrand of interest.
@@
 The dimension $s$ need not equal $d$.
-We will use RQMC points $\bm{z}\in[0,1]^{s+1}$.
-We partition $\bm{z}$ into its first component $v$ that we use to select a stratum $\ell$ and the subsequent $s$ components $\bm{u}$ that we use to sample $\bm{x}\sim P_{\ell}$.
-That is, $\bm{z}=(v,u_{1},u_{2},\dots,u_{s})$.
-For a function $\bm{\ell}:[0,1]\to\{1,2,\dots,L\}$, we will sample the $i$’th point using sample stratum $\bm{\ell}(v_{i})$.
-The pre-image of each $\ell\in\{1,2,\dots,L\}$ under $\bm{\ell}$ will always be a sub-interval of $[0,1]$.
-We let $\beta_{\ell}$ be the length of that subinterval.
-We abbreviate $\bm{\ell}(v_{i})$ to $\ell(i)$ in some expressions.
-Having selected stratum $\ell(i)$, we draw $\bm{x}_{i}\in^{d}$ via $\bm{x}_{i}=\phi_{\ell(i)}(\bm{u}_{i})$.
+[…]
 We now estimate $\mu={\mathbb{E}}(g(\bm{x}))$ by
@@
 $\mu=\int_{[0,1]^{s+1}}f(\bm{z})\mathrm{\,d}\bm{z}.$
-We also define the within-stratum integrands individually as functions of $[0,1]^{s}$. The within-stratum integrands and integrals are
+Because each RQMC point $\bm{z}_{i}$ is uniformly distributed on $0,1]^{s+1}$, the estimator $\hat{\mu}$ is unbiased for $\mu$.
+We also define the within-stratum integrands individually as functions of $0,1]^{s}$. The within-stratum integrands and integrals are
 $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))\quad\text{and}\quad\mu_{\ell}=\int_{[0,1]^{s}}h_{\ell}(\bm{u})\mathrm{\,d}\bm{u}$
@@
 Since we already have a dependence on $\ell$ in $h_{\ell}$ we could also replace $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))$ by $h_{\ell}(\bm{u})=g_{\ell}(\phi_{\ell}(\bm{u}))=g_{\ell}(\bm{x})$ to allow the function being applied to $\bm{x}$ to vary with $\ell$.
+That is $\mu=\sum_{\ell=1}^{L}\alpha_{\ell}{\mathbb{E}}(g_{\ell}(\bm{x})\!\mid\!\bm{x}\sim P_{\ell})$.
 2.2 Sampling the strata
@@
 We will always order the strata so that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$.
-That brings an advantage in Section 2.5.
+This ordering brings a sampling advantage in Section 2.5.
+It is ordinarily the same ordering as the $\alpha_{\ell}$, though we consider exceptions at the end of Section 3.1.
 We will assume that the sampling values $v_{1},\dots,v_{n}$ that we use are ‘stratified’.
 By this we mean that every interval $I_{k}=[(k-1)/n,k/n)$ contains precisely one of $v_{1},\dots,v_{n}$ for $k=1,\dots,n$.
-The next proposition shows that if every $\beta_{\ell}>2/n$, then every $n_{\ell}>0$ when $v_{i}$ are stratified.
+The next proposition shows that $n_{\ell}>0$ whenever $\beta_{\ell}>2/n$.
 Proposition 1.
-For $n\geqslant 1$, let $v_{1},\dots,v_{n}$ be stratified. For $[A,B)\subset[0,1)$, let $n_{*}$ be the number of $v_{i}\in[A,B)$. Then
+For $n\geqslant 1$, let $v_{1},\dots,v_{n}$ be stratified. For $A,B)\subset[0,1)$, let $n_{*}$ be the number of $v_{i}\in[A,B)$. Then
 $\lceil n\beta\rceil-2\leqslant n_{*}\leqslant\lfloor n\beta\rfloor+2$ (5)
@@
 Proof.
-Write $A=r/n+a$ and $B=s/n+b$ for integers $0\leqslant r\leqslant s<n$ and $a,b\in[0,1/n)$.
-Then $\beta=t/n+c$ for $t=s-r$ and $-1/n<c<1/n$.
-If $r=s$, then $0\leqslant n_{*}\leqslant 1$ and then (6) holds trivially.
-For $s>r$ we may write
-$[A,B)=[A,(r+1)/n)\,\cup\,[(r+1)/n,s/n)\,\cup\,[s/n,s/n+b).$
-Then $[A,B)$ contains at least $s-r-1$ of the intervals $I_{k}=[(k-1)/n,k/n)$ for $k=1,\dots,n$ and it intersects with at most $s-r+1$ of them.
-Therefore $|n_{*}-(s-r)|\leqslant 1$.
-Now $\lceil n\beta\rceil=s-r+\lceil nc\rceil$ and so $s-r\geqslant\lceil n\beta\rceil-1$ establishing the lower limit in (6).
-Similarly $\lfloor n\beta\rfloor=s-r+\lfloor nc\rfloor$ so $s-r\leqslant\lfloor n\beta\rfloor+1$ establishing the upper limit in (6).
-∎
+[…]
 2.3 Local and star discrepancy
-For $\bm{a}\in[0,1)^{d}$, the anchored box $[\bm{0},\bm{a})$ has volume $\mathrm{Vol}([\bm{0},\bm{a}))=\prod_{j=1}^{d}a_{j}$.
+For $\bm{a}\in[0,1)^{d}$, the anchored box $\bm{0},\bm{a})$ has volume $\mathrm{Vol}([\bm{0},\bm{a}))=\prod_{j=1}^{d}a_{j}$.
 The local discrepancy of $n$ points $\bm{x}_{1},\dots,\bm{x}_{n}\in[0,1)^{d}$ at $\bm{a}$ is
@@
 $D_{n}^{*}=D_{n}^{*}(\bm{x}_{1},\dots,\bm{x}_{n})=\sup_{\bm{a}\in[0,1)^{d}}|\delta(\bm{a})|.$
-The Koksma-Hlawka inequality (Hickernell, 2014) is that
+The Koksma-Hlawka inequality (Hickernell 2014) is that
 $\Bigl|\frac{1}{n}\sum_{i=1}^{n}f(\bm{x}_{i})-\int_{[0,1]^{d}}f(\bm{x})\mathrm{\,d}\bm{x}\Bigr|\leqslant D_{n}^{*}(\bm{x}_{1},\dots,\bm{x}_{n})V_{{\mathrm{HK}}}(f)$ (6)
@@
 Here we present the defining properties of digital nets.
-They exist in integer bases $b\geqslant 2$ (Dick and Pillichshammer, 2010) but we only consider base $2$ here.
+They exist in integer bases $b\geqslant 2$ (Dick & Pillichshammer 2010) but we only consider base $2$ here.
 For a dimension $d\geqslant 1$, elementary intervals in base $2$ are dyadic hyperrectangles of the form
 $H(\bm{c},\bm{k})=\prod_{j=1}^{d}\Bigl[\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr)$
-for integers $k_{j}\geqslant 0$ and $0\leqslant c_{j}<2^{k_{j}}$. We will use $|\bm{k}|=\sum_{j=1}^{d}k_{j}$, so $H(\bm{c},\bm{k})$ has volume $2^{-|\bm{k}|}$.
+for integers $k_{j}\geqslant 0$ and $0\leqslant c_{j}<2^{k_{j}}$. We use $|\bm{k}|=\sum_{j=1}^{d}k_{j}$, so $H(\bm{c},\bm{k})$ has volume $2^{-|\bm{k}|}$.
 For integers $m\geqslant t\geqslant 0$ the $n=b^{m}$ points $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1)^{d}$ are a $(t,m,d)$-net in base $b$ if $H(\bm{c},\bm{k})$ contains precisely $nb^{-|\bm{k}|}=b^{m-|\bm{k}|}$ of them, whenever $m-|\bm{k}|\geqslant t$.
 Smaller $t$ are better.
-We always have $t\leqslant m$ because $t=m$ just means that all the points are in $[0,1)^{d}$.
+We always have $t\leqslant m$ because $t=m$ just means that all the points are in $0,1)^{d}$.
 A $(t,d)$-sequence in base $2$ is an infinite sequence of values $\bm{x}_{i}\in[0,1]^{d}$, with the following property: any subsequence $\bm{x}_{a2^{r}+1},\dots,\bm{x}_{(a+1)2^{r}}$ for integers $a\geqslant 0$ and $r\geqslant m$ is a $(t,r,d)$-sequence in base $2$.
@@
 Consecutive pairs of those $(t,m,d)$-nets yield an infinite sequence of $(t,m+1,d)$-nets of which pairs form $(t,m+2,d)$-nets and so on without limit.
-We will use digital nets in base $2$ taken from Sobol’ sequences using the direction numbers from Joe and Kuo (2008).
+We will use digital nets in base $2$ taken from Sobol’ sequences using the direction numbers from Joe & Kuo (2008).
 We call a $(t,m,d)$-net in base $2$ stratified if for each component $j=1,\dots,d$, the values $x_{1j},\dots,x_{nj}$ are stratified.
@@
 The stratification property means we could use the $j$’th variable in our $(t,m,d)$-net to select $\ell(i)$ for any $j$.
-Stratification implies the ‘projection regularity’ property of Sloan and Joe (1994) in which we never get $x_{ij}=x_{i^{\prime}j}$ for any $i\neq i^{\prime}$ and any $j$.
+Stratification implies the ‘projection regularity’ property of Sloan & Joe (1994) in which we never get $x_{ij}=x_{i^{\prime}j}$ for any $i\neq i^{\prime}$ and any $j$.
 The Sobol’ points we use are ones that have been scrambled.
-The scrambles we consider are either nested uniform scrambles (Owen, 1995) or random linear scrambles with digital shifts (Matoušek, 1998).
+The scrambles we consider are either nested uniform scrambles (Owen 1995) or random linear scrambles with digital shifts (Matoušek 1998).
 Scrambling makes each $\bm{z}_{i}\sim\mathbf{U}[0,1]^{d}$ individually while $\bm{z}_{1},\dots,\bm{z}_{n}$ remain a $(t,m,d)$-net collectively.
 We will estimate $\mu$ by $\hat{\mu}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are a scrambled $(t,m,s+1)$-net in base $2$.
-This scrambled net integration has several desirable properties listed in Owen and Rudolf (2020):
+This scrambled net integration has several desirable properties listed in Owen & Rudolf (2020):
   1. 1)
@@
 The scrambled Sobol’ nets are stratified.
+Digital nets in base $2$ have sample sizes $n=2^{m}$ by definition.
+Arbitrary sample sizes $n$ are available as the first $n$ points of scrambled Sobol’ sequences.
+Property $1$ holds for any finite $n$ there and Property 2 holds for any sequence $n\to\infty$ (Owen & Rudolf 2020).
+The other properties are specific to $n=2^{m}$.
 From point 3 above we know that ${\mathrm{Var}}(\hat{\mu})=o(n^{-1})$ for $f\in L_{2}$.
-The Sobol’ nets have $D_{n}^{*}=O(n^{-1}\log(n)^{d-1})$ (Niederreiter, 1992, Theorem 4.10).
+The Sobol’ nets have $D_{n}^{*}=O(n^{-1}\log(n)^{d-1})$ (Niederreiter 1992, Theorem 4.10).
 Scrambled nets are still nets, so this holds for them too.
-Then by the Koksma-Hlawka inequality, ${\mathrm{Var}}(\hat{\mu})=O(n^{-2+\epsilon})$ holds for any $\epsilon>0$ when $V_{{\mathrm{HK}}}(f)<\infty$, which we describe as $f$ being in BVHK.
+Then by the Koksma-Hlawka inequality, ${\mathrm{Var}}(\hat{\mu})=O(n^{-2+\epsilon})$ holds for any $\epsilon>0$ when $V_{{\mathrm{HK}}}(f)<\infty$.
+Then $f$ is of bounded variation in the sense of Hardy and Krause, which we describe as $f$ being in BVHK.
 For regular $f$ described in point 4 above we have ${\mathrm{Var}}(\hat{\mu})=O(n^{-3+\epsilon})$.
-Higher order digital nets can attain even better convergence rates under correspondingly stronger regularity (Dick, 2011), though they are not widely used for higher dimensions.
+Higher order digital nets can attain even better convergence rates under correspondingly stronger regularity (Dick 2011), though they are not widely used for higher dimensions.
 It is common to write these powers of $\log(n)$ as $O(n^{\epsilon})$.
-For a discussion of that point, see Owen and Pan (2022).
-While convergence $O(n^{-\rho})$ for $\rho\in\{1,2,3\}$ are widely studied, these come from error bounds and the actual convergence rate need not be near an integer.
+For a discussion of that point, see Owen & Pan (2022).
+While variance rates $O(n^{-\rho})$ for $\rho\in\{1,2,3\}$ are widely studied, these come from error bounds.
+The actual convergence rate need not be near an integer.
 Liu (2025) shows examples where certain singular behavior of the integrand can give any rate between $1$ and $2$.
@@
 Proof.
-For $S_{\ell}$ defined at (4), sample $i\in S_{\ell}$ holds if and only if $B_{\ell-1}\leqslant z_{i1}<B_{\ell}$ or equivalently
-$\sum_{1\leqslant r<\ell}2^{-\kappa_{r}}\leqslant z_{i1}<\sum_{1\leqslant r\leqslant\ell}2^{-\kappa_{r}}$
-which we can write as
-$\frac{\gamma}{2^{\kappa_{\ell}}}\leqslant z_{i1}<\frac{\gamma+1}{2^{\kappa_{\ell}}}$ (8)
-for integer $\gamma=\sum_{1\leqslant r<\ell}2^{\kappa_{\ell}-\kappa_{r}}\geqslant 0$.
-Because the net is stratified, the number of observations $i$ that satisfy equation (8) is $2^{m-\kappa_{\ell}}=2^{m_{\ell}}$.
-Then $\{u_{i}\mid i\in S_{\ell}\}\subset[0,1)^{s}$ has cardinality $2^{m_{\ell}}$ making it a $(t_{\ell},m_{\ell},s)$-net in base $2$ for some value $t_{\ell}\leqslant m_{\ell}$.
-Now consider a dyadic box in $[0,1]^{s}$ of the form
-$\prod_{j=1}^{s}\Bigl{[}\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr{)}.$
-This box has volume $2^{-|\bm{k}|}$. The number of points $\bm{u}_{i}$ from stratum $\ell$ in this box equals the number of points $\bm{z}_{i}\in[0,1)^{s+1}$ that belong to
-$\Bigl{[}\frac{\gamma}{2^{\kappa_{\ell}}},\frac{\gamma+1}{2^{\kappa_{\ell}}}\Bigr{)}\times\prod_{j=1}^{s}\Bigl{[}\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr{)}.$
-If $m-\kappa_{\ell}-|\bm{k}|\geqslant t$, then by the digital net property of $\bm{z}_{1},\dots,\bm{z}_{n}$, this number is $2^{m-\kappa_{\ell}-|\bm{k}|}$.
-That is, if $m_{\ell}\geqslant|\bm{k}|+t$, then the number of such points is $2^{m_{\ell}-|\bm{k}|}$.
-∎
+[…]
 To motivate the constraint that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$, consider what would happen if we took $\beta_{1}=1/8$, $\beta_{2}=1/2$, $\beta_{3}=1/4$ and $\beta_{4}=1/8$.
@@
 Proof.
-The local discrepancy of $\bm{z}_{1},\dots,\bm{z}_{n}$ at the point $(b,\bm{c})\in[0,1]^{s+1}$ is
-$\delta(b,\bm{c})=\frac{1}{n}\sum_{i=1}^{n}\bm{1}\{0\leqslant v_{i}<b\}\bm{1}\{\bm{z}_{i}\in[\bm{0},\bm{c})\}-b\times\mathrm{Vol}([\bm{0},\bm{c}))$
-where $\mathrm{Vol}([\bm{0},\bm{c}))=\prod_{j=1}^{s}c_{j}$. Because $\bm{z}_{1},\dots,\bm{z}_{n}$ are a digital net, $|\delta(b,\bm{c})|=O(n^{-1}\log(n)^{s})$.
-The sample points for stratum $\ell$ are the $n_{\ell}$ points $\bm{x}_{i}$ for which $B_{\ell-1}\leqslant v_{i}<B_{\ell}$.
-By equation (6), $n_{\ell}=n\beta_{\ell}+O(1)$.
-The local discrepancy within stratum $\ell$ at $\bm{c}\in[0,1)^{s}$ is
-$\ \quad\frac{1}{n_{\ell}}\sum_{i=1}^{n}\bm{1}\{B_{\ell-1}\leqslant v_{i}<B_{\ell}\}\bm{1}\{\bm{x}_{i}\in[\bm{0},\bm{c})\}-\mathrm{Vol}([\bm{0},\bm{c}))$
-$=\frac{n}{n_{\ell}}\Bigl{(}\beta_{\ell}\mathrm{Vol}([\bm{0},\bm{c}))+\delta(B_{\ell},\bm{c})-\delta(B_{\ell-1},\bm{c})\Bigr{)}-\mathrm{Vol}([\bm{0},\bm{c})).$
-Its absolute value is at most
-$\Bigl{|}\frac{n\beta_{\ell}}{n}-1\Bigr{|}\mathrm{Vol}([\bm{0},\bm{c}))+\frac{2n}{n_{\ell}}O(n^{-1}\log(n)^{s})$ $=O(n_{\ell}^{-1}\log(n_{\ell}/\beta_{\ell})^{s})$
-$=O(n_{\ell}^{-1}\log(n_{\ell})^{s})$
-and so the star discrepancy of $\bm{u}_{i}$ for $i\in S_{\ell}$ is $O(n_{\ell}^{-1}\log(n_{\ell})^{s})$. ∎
+[…]
 3 Choosing $\beta_{\ell}$ for given $\alpha_{\ell}$
@@
 We will show that choosing $\beta_{\ell}$ based on the wrong value of $\rho$ can adversely affect the implied constant in the asymptotic MSE while leaving the convergence rate unchanged.
-Values of $\rho$ equal to $1$, $2$ and $3$ correspond to upper bounds on the convergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases.
+Values of $\rho$ equal to $1$, 22 and $3$ correspond to upper bounds on the convergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases.
 It often happens that the empirical MSE of an RQMC rule follows an apparent rate of $n^{-\rho}$ for a value of $\rho$ that is not close to an integer.
 Unbounded integrands are not in BVHK, and yet they will have an MSE that is $o(n^{-1})$.
-Those often appear to have $21<\rho<$ 2.
+Those often appear to have $1<\rho<2$.
 Our sampling algorithm will deliver $n_{\ell}$ with $|n_{\ell}-n\beta_{\ell}|\leqslant 2$ by Proposition 1.
@@
 For $\rho>2$, we have the reverse situation.
-Changing $n_{\ell}$ by even ±$\pm 1$ can worsen the convergence rate because changing $\hat{\mu}$ by $O(1/n)$ is a bigger change than the RMSE of $O(n^{-\rho/2})$.
+Changing $n_{\ell}$ by even $\pm 1$ can worsen the convergence rate because changing $\hat{\mu}$ by $O(1/n)$ is a bigger change than the RMSE of $O(n^{-\rho/2})$.
 A convergence rate with $\rho>2$ for scrambled Sobol’ points is only available for sample sizes that are powers of two.
-Section 4 shows how to obtain $n_{\ell}$ equal to powers of two.
+Section 4 shows how to enforce that all $n_{\ell}$ are equal to powers of two.
 Either way, we can assume that $n_{\ell}=n\beta_{\ell}$. Then with $\omega_{\ell}=\alpha_{\ell}/\beta_{\ell}$, we can rewrite (4) as
@@
 3.1 Optimal sample sizes
-Under the model (9),
-${\mathrm{Var}}(\hat{\mu})={\mathrm{Var}}\biggl{(}\sum_{\ell=1}^{L}\alpha_{\ell}\hat{\mu}_{\ell}\biggr{)}\leqslant\biggl{(}\sum_{\ell=1}^{L}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}\biggr{)}^{2}$ (10)
-where the upper bound is from the very conservative assumption that the estimates within all $L$ strata have unit correlations.
-With no prior information on the relative sizes of $\tau_{\ell}$ we adopt
-$C_{1}(\rho)=C_{1}(\rho;n_{1},\dots,n_{L})=\sum_{\ell=1}^{L}\alpha_{\ell}n_{\ell}^{-\rho/2}$ (11)
-as a criterion to minimize subject to $\sum_{\ell=1}^{L}n_{\ell}=n$. We form the Lagrangian
-$G_{1}=\sum_{\ell=1}^{L}\alpha_{\ell}n_{\ell}^{-\rho/2}+\lambda\sum_{\ell=1}^{L}n_{\ell}$
-and setting $\partial G_{1}/\partial n_{\ell}$ to zero yields a solution with
-$n_{\ell}\propto\alpha_{\ell}^{2/(\rho+2)}.$ (12)
-We then suppose that the user chooses integers $n_{\ell}$ summing to $n=2^{m}$ that are nearly proportional to $\alpha_{\ell}^{2/(\rho+2)}$.
-RQMC points have a space filling property.
-By construction, the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$.
-We might then anticipate primarily non-positive correlations among the $\hat{\mu}_{\ell}$.
-This motivates a different criterion
-$C_{0}(\rho)=C_{0}(\rho;n_{1},\dots,n_{L})=\sum_{\ell=1}^{L}\alpha_{\ell}^{2}n_{\ell}^{-\rho}$ (13)
-based on zero correlations among the $\hat{\mu}_{\ell}$. Forming the Lagrangian for this criterion and setting to zero the partial derivatives with respect to $n_{\ell}$ leads to
+Under the model (7), ${\mathrm{Var}}(\hat{\mu})={\mathrm{Var}}\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}\hat{\mu}_{\ell}\bigr)\leqslant\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}\bigr)^{2}$where the upper bound is from the very conservative assumption that the estimates within all $L$ strata have unit correlations.
+RQMC points have a space filling property: the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$.
+Then a zero or even negative correlation among stratum means is more plausible.
+Under a working model of 0 inter-stratum correlation, ${\mathrm{Var}}(\hat{\mu})\approx\sum_{\ell=1}^{L}\alpha_{\ell}^{2}\tau_{\ell}/n_{\ell}^{\rho}$.
+Then with no prior information on the relative sizes of $\tau_{\ell}$ we adopt the criterion
+$C_{0}(\rho)=C_{0}(\rho;n_{1},\dots,n_{L})=\sum_{\ell=1}^{L}\alpha_{\ell}^{2}n_{\ell}^{-\rho}.$ (8)
+The conservative criterion assuming unit correlations is $C_{1}(\rho)=\sum_{\ell=1}^{L}\alpha_{\ell}n_{\ell}^{-\rho/2}$.
+Forming the Lagrangian for (8) with the constraint $\sum_{\ell}n_{\ell}=n$ and setting to zero the partial derivatives with respect to $n_{\ell}$ leads to
 $n_{\ell}\propto\alpha_{\ell}^{2/(\rho+1)}.$ (9)
-In both criteria (12) and (14), higher convergence rates $\rho$ tend to equalize the $n_{\ell}$.
+We then suppose that the user chooses integers $n_{\ell}$ summing to $n=2^{m}$ that are nearly proportional to $\alpha_{\ell}^{2/(\rho+1)}$.
+The conservative criterion 1 leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho+2)}$.
+For both criteria higher convergence rates $\rho$ tend to equalize the $n_{\ell}$.
 We can think of $n_{\ell}=n/L$ as arising in the limit as $\rho\to\infty$.
 We will also see equal $n_{\ell}$ in a minimax argument in Section 5.
-We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for ansatzes 00 and $1$, and three different hypotheses on $\rho$.
+The conservative criterion 1 favors stratum sizes more nearly equal than under criterion 0 while not exactly equal as the even more conservative minimax argument gives.
+We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for criteria 0 and $1$, and three different hypotheses on $\rho$.
 They lead to the four different rates given in Table 1.
-The case for ansatz 00 and $\rho=1$ (uncorrelated estimates at the MC rate) is the only one with a proportional allocation.
+The case for criterion 0 and $\rho=1$ (uncorrelated estimates at the MC rate) is the only one with a proportional allocation.
 The other choices all allocate more than a proportional number of samples to the strata with smaller $\alpha_{\ell}$.
-This is also true for the plain MC rate ($\rho=1$) under the conservative ansatz 1.
-When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then equations (12) and (14) do not distinguish between different values of $\rho$.
+This is also true for the plain MC rate ($\rho=1$) under the conservative criterion 1.
+When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then our criteria do not distinguish between different values of $\rho$.
 After raising $\alpha_{\ell}$ to any power they are still equal and then a near equal integer allocation is appropriate.
@@
 In that case, the proportionality for $C_{0}$ and $\rho=1$ matches the Neyman allocation from survey sampling, in Neyman (1934).
-When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$, then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}/c_{\ell}$ in Table 1.
+When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$, then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}c_{\ell}^{-1/2}$ in Table 1.
+When using this extra information, we should order the strata to be in non-increasing order of $\alpha_{\ell}\tau_{\ell}^{1/2}c_{\ell}^{-1/2}$ for either criterion.
 $\rho:$ 1 2 3
@@
 $C_{1}$ $\alpha_{\ell}^{2/3}$ $\alpha_{\ell}^{1/2}$ $\alpha_{\ell}^{2/5}$
-Table 1: This table shows ideal rates for $n_{\ell}$ as functions of $\alpha_{\ell}$.
-Under ansatz 00, estimates from different strata are uncorrelated.
-Ansatz $1$ uses a bound assuming that those estimates are perfectly correlated.
-These rates apply when the variance within stratum $\ell$ decreases proportionally to $n_{\ell}^{-\rho}$.
-If $\rho$ varies from stratum to stratum, taking the value $\rho_{\ell}$ in stratum $\ell$, then the Lagrange multiplier argument leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho_{\ell}+1)}$.
-Strata with higher $\rho_{\ell}$ get larger $n_{\ell}$ and $\beta_{\ell}=n_{\ell}/n$ does not depend on $n$.
+Table 1: This table shows ideal rates for $n_{\ell}$ as functions of $\alpha_{\ell}$ with $C_{0}$ and $C_{1}$ for criteria 0 and 11 respectively.
+These rates assume the variance within stratum $\ell$ is proportional to $n_{\ell}^{-\rho}$.
+[…]
 3.2 Suboptimality
@@
 We measure inefficiency by the ratio of the MSE we get if we design with a value $\gamma$ when we should have used $\rho$, so smaller values are better, and $1$ is best possible.
-In ansatz 0, we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
+Under criterion 0, we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
 $I_{0}(\gamma\!\mid\!\rho)$ $=\frac{\sum_{\ell=1}^{L}\alpha^{2}_{\ell}\Bigl(n\alpha_{\ell}^{2/(\gamma+1)}\bigm/\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+1)}\Bigr)^{-\rho}}{\sum_{\ell=1}^{L}\alpha^{2}_{\ell}\Bigl(n\alpha_{\ell}^{2/(\rho+1)}\bigm/\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)}\Bigr)^{-\rho}}$
@@
 Only the lead constant is affected.
-For ansatz 1, assuming a rate $n_{\ell}^{-\gamma}$ instead of $n_{\ell}^{-\rho}$ would lead to $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+2)}$.
-To put the inefficiency in squared error terms, comparable to $I_{0}$ we use $C^{2}_{1}(\gamma)/C^{2}_{1}(\rho)$, getting
-$I_{1}(\gamma\!\mid\!\rho)$ $=\left(\frac{\sum_{\ell=1}^{L}\alpha_{\ell}\bigl{(}n\alpha_{\ell}^{2/(\gamma+2)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+2)}\bigr{)}^{-\rho/2}}{\sum_{\ell=1}^{L}\alpha_{\ell}\bigl{(}n\alpha_{\ell}^{2/(\rho+2)}\bigm{/}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}\bigr{)}^{-\rho/2}}\right)^{2}$
-$=\frac{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{1-\rho/(\gamma+2)}\bigr{)}^{2}\bigl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+2)}\bigr{)}^{\rho}}{\bigl{(}\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}\bigr{)}^{\rho+2}}.$
-We think that ansatz 0 makes a more reasonable assumption than ansatz 1. Then a cautious choice is to take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma_{0}+1)}$ where
+The analogous formula for criterion 1 is
+$I_{1}(\gamma\!\mid\!\rho)$ $=\frac{\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}^{1-\rho/(\gamma+2)}\bigr)^{2}\bigl(\sum_{k=1}^{L}\alpha_{k}^{2/(\gamma+2)}\bigr)^{\rho}}{\bigl(\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}\bigr)^{\rho+2}}.$
+Under criterion 0 a cautious choice is to take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma_{0}+1)}$ where
 $\gamma_{0}=\operatorname*{argmin}_{1\leqslant\gamma\leqslant 3}\max_{1\leqslant\rho\leqslant 3}I_{0}(\gamma\!\mid\!\rho).$ (11)
@@
 For $I_{1}$ we want $1-\rho/(\gamma+2)\geqslant 0$ or $\gamma\geqslant\rho-2$.
-Since we already had $\gamma\geqslant\rho-1$ using ansatz 0, this is not a further constraint.
+Since we already had $\gamma\geqslant\rho-1$ using criterion 0, this is not a further constraint.
 When we know that $\rho\leqslant 2$, then any $\gamma\geqslant 1$ avoids a negative exponent for $\alpha_{\ell}$.
@@
 Proof.
-See Appendix A. ∎
+[…]
 If the unknown true rate is $\rho=2$, we are better off choosing $\gamma=3$ than $\gamma=1$. Proposition 5 has this ‘bet on optimism’ result.
@@
 Proof.
-From equation (3.2)
-$I_{0}(1\!\mid\!2)=\frac{L}{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{2/3}\bigr{)}^{3}}\quad\text{and}\quad I_{0}(3\!\mid\!2)=\frac{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{1/2}\bigr{)}^{2}}{\bigl{(}\sum_{\ell=1}^{L}\alpha_{\ell}^{2/3}\bigr{)}^{3}}.$
-These expressions have the same denominator and then the result follows from the Cauchy-Schwarz inequality applied to the numerators. ∎
+[…]
 4 Making each $n_{\ell}$ a power of two
@@
 The advantage of having every $n_{\ell}$ be a power of two is that we then get scrambled net sampling within each of the $L$ strata as shown by Proposition 2.
-Sample sizes that are a power of two are a best practice in scrambled Sobol’ sampling (Owen, 2020).
+Sample sizes that are a power of two are a best practice in scrambled Sobol’ sampling (Owen 2020).
 The results with MSE $O(n^{-3+\epsilon})$ for scrambled Sobol’ points require such sample sizes.
@@
 Figure 1: These are the partitions of unity into $L$ negative powers of $2$ for $L\in\{4,5,6,7,8\}$.
-The number of such partitions as a function of $L$ is given by sequence A002572 at the online encyclopedia of integer sequences, (OEIS Foundation Inc., 2023) along with motivating applications such as counting the number of rooted binary trees or, more recently, counting nonequivalent compact Huffman codes (Elsholtz et al.
+The number of such partitions as a function of $L$ is given by sequence A002572 at the online encyclopedia of integer sequences, (OEIS Foundation Inc.
+2023) along with motivating applications such as counting the number of rooted binary trees or, more recently, counting nonequivalent compact Huffman codes (Elsholtz et al.
 (2024)).
@@
 Proof.
-First, if $n^{\prime}>0$ then we can show that there is an ‘eligible’ index $\ell$ with $n_{\ell}\leqslant n^{\prime}$, that can be doubled without causing $\sum_{\ell=1}^{L}n_{\ell}>n$.
-At any stage in the algorithm, $n_{\ell^{*}}$ for $\ell^{*}=\operatorname*{argmin}_{1\leqslant\ell\leqslant L}n_{\ell}$ is a power of two.
-It divides $n$ and it divides every $n_{\ell}$, so it also divides the nonzero value $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$, hence $n_{\ell^{*}}\leqslant n^{\prime}$.
-Finally, every step with $n^{\prime}>0$ reduces $n^{\prime}$, and so $n^{\prime}$ must reach zero in a finite number of steps starting from $n^{\prime}=n-L$.
-∎
-Under ansatz 0, the ideal proportion of samples in stratum $\ell$ is
+[…]
+Under criterion 0, the ideal proportion of samples in stratum $\ell$ is
 $\xi_{\ell,\rho}=\alpha_{\ell}^{2/(\rho+1)}\Bigm/\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)}.$
@@
 $\ell^{*}_{0}=\operatorname*{argmax}_{1\leqslant\ell\leqslant L}\frac{\xi_{\ell,\rho}}{2^{m_{\ell}}}\bm{1}\Bigl\{2^{m_{\ell}}\leqslant n-\sum_{j=1}^{L}2^{m_{j}}\Bigr\}$
-which we consider to be the most ‘under-represented’ of the eligible strata. The analogous algorithm under ansatz $1$ is to double the sample size of stratum
+which we consider to be the most ‘under-represented’ of the eligible strata. The analogous algorithm under criterion $1$ is to double the sample size of stratum
 $\ell^{*}_{1}=\operatorname*{argmax}_{1\leqslant\ell\leqslant L}\frac{\xi_{\ell,\rho}}{2^{m_{\ell}}}\bm{1}\Bigl\{2^{m_{\ell}}\leqslant n-\sum_{j=1}^{L}2^{m_{j}}\Bigr\}$
@@
 $\xi_{\ell,\rho}=\alpha_{\ell}^{2/(\rho+2)}\Bigm/\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+2)}.$
-The pseudo-code for forward allocation algorithm for ansatz 0 is given in Algorithm 1.
+The pseudo-code for forward allocation algorithm for criterion 0 is given in Algorithm 1.
 Algorithm 1 Forward Stratified Allocation
@@
 Here we show a minimax property of near-equal allocation. We do not assume equality among the $\tau_{\ell}$. For any integer $N\geqslant L$, let
-$\Delta_{N}$ =$=\biggl{\{}\bm{n}=(n_{1},\dots,n_{L})\in\mathbb{N}^{L}\Bigm{|}\sum_{\ell=1}^{L}n_{\ell}=N\biggr{\}},\quad\text{and}\quad$
-$\mathcal{T}$ =$=\Bigl{\{}\bm{\tau}=(\tau_{1},\dots,\tau_{L})\in[0,\infty)^{L}\bigm{|}\max_{1\leqslant\ell\leqslant L}\tau_{\ell}>0\Bigr{\}}$
+$\Delta_{N}$ $=\biggl\{\bm{n}=(n_{1},\dots,n_{L})\in\mathbb{N}^{L}\Bigm|\sum_{\ell=1}^{L}n_{\ell}=N\biggr\},\quad\text{and}\quad$
+$\mathcal{T}$ $=\Bigl\{\bm{\tau}=(\tau_{1},\dots,\tau_{L})\in[0,\infty)^{L}\bigm|\max_{1\leqslant\ell\leqslant L}\tau_{\ell}>0\Bigr\}$
 be spaces of feasible sample sizes and possible variance constants, respectively. We allow some $\tau_{\ell}=0$ but there must be at least one $\tau_{\ell}>0$ to make the problem interesting.
 We consider the increase in variance that we get using sample sizes $\bm{n}$ compared to some better sample sizes $\tilde{\bm{n}}$ that we might otherwise have used.
-For ansatzes 00 and $1$ we define suboptimality ratios
+For criteria 0 and $1$ we define suboptimality ratios
 $R_{0}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}n_{\ell}^{-\rho}}{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}\tilde{n}_{\ell}^{-\rho}}\quad\text{and}\quad R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}}{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}\tilde{n}_{\ell}^{-\rho/2}}$
@@
 Proposition 7.
-For both ansatzes $a\in\{0,1\}$,
+For both criteria $a\in\{0,1\}$,
 $\min_{\bm{n}\in\Delta_{N}}\max_{\tilde{\bm{n}}\in\Delta_{N}}\max_{\bm{\tau}\in\mathcal{T}}R_{a}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau})=R_{a}(\bm{n}^{*}\!\mid\!\tilde{\bm{n}}^{*};\bm{\tau}^{*})$ (13)
@@
 Proof.
-The same argument applies to both ansatzes.
-We prove it for ansatz 1.
-Fixing $\bm{n}$ and $\tilde{\bm{n}}$ we see that the maximum over $\bm{\tau}$ of $R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau})$ arises when $\tau_{\ell}=0$ for all $\ell$ except $\ell^{*}(\bm{n},\tilde{\bm{n}})=\operatorname*{argmax}_{\ell}(\tilde{n}_{\ell}/n_{\ell})^{\rho/2}$ with $\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}$ then taking any value $\lambda>0$.
-Now, we can go back to the initial problem and rewrite (18) as:
-$\min_{\bm{n}\in\Delta_{N}}\max_{\tilde{\bm{n}}\in\Delta_{N}}\Bigl{(}\frac{\tilde{n}_{\ell^{*}(\bm{n},\tilde{\bm{n}})}}{{n}_{\ell^{*}(\bm{n},\tilde{\bm{n}})}}\Bigr{)}^{\rho/2}$
-after noting that the factor $\alpha_{\ell^{*}(\bm{n},\tilde{\bm{n}})}\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}^{1/2}$ cancels from numerator and denominator.
-To maximize this ratio over $\tilde{\bm{n}}$ one takes $\tilde{\bm{n}}$ to be the vector with value $N-L+1$ in component $\ell^{*}=\operatorname*{argmin}_{\ell}n_{\ell}$ and ones elsewhere.
-Then the minimax problem becomes $\min_{\bm{n}\in\Delta_{N}}\bigl{(}(N-L+1)/\min_{1\leqslant\ell\leqslant N}n_{\ell}\bigr{)}^{\rho/2}$.
-The solution is to maximize the minimum value of $n_{\ell}$.
-We do that by making $\min_{\ell}n_{\ell}=\lfloor N/L\rfloor$.
-∎
+[…]
 When we must have $n_{\ell}=2^{m_{\ell}}$ for integers $m_{\ell}\geqslant 0$, there is a corresponding minimax result that takes the $n_{\ell}$ as nearly equal as possible.
@@
   2. 2)
-    A plain RQMC estimator: $\hat{\mu}_{\mathrm{RQMC}}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are the first $n$ points of a scrambled Sobol’ sequence in $[0,1]^{2}$.
+    A plain RQMC estimator: $\hat{\mu}_{\mathrm{RQMC}}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are the first $n$ points of a scrambled Sobol’ sequence in $0,1]^{2}$.
   3. 3)
@@
 For each estimator we compute $R=500$ replicates at sample sizes $n=2^{m}$ for $3\leqslant m\leqslant 12$ and then graph the sample variances.
-We used nested uniform scrambling (Owen, 1995) to do the scrambling.
+We used nested uniform scrambling (Owen 1995) to do the scrambling.
 The results are shown in Figure 2.
@@
 Figure 2: Variance of the 5 estimators for the estimation of ${\mathbb{E}}(f(\bm{z}))$ versus the sample size $n$, for the toy integrand.
+The top to bottom ordering of the curves in the legend is the same as they have at $n=4096$.
 Four of the curves are nearly parallel to reference lines of a given slope.
@@
 Great uncertainty often arises from weather variables or customer behavior.
-Our example is the Saint-Venant flood model from Limbourg and De Rocquigny (2010).
+Our example is the Saint-Venant flood model from Limbourg & De Rocquigny (2010).
 A more precise flood model would be based on Navier-Stokes equations that have to be solved numerically.
@@
 $K_{s}$ Gamma (shape $90$, scale $1/3$) Gamma (shape $15$, scale $1$)
-$Z_{v}$ Uniform ($49$,$51$) none
-$Z_{m}$ Uniform ($54$,$56$) none
+$Z_{v}$ Uniform (4949,$51$) none
+$Z_{m}$ Uniform (5454,$56$) none
 Table 2: Here are nominal distributions for four variables in the Saint-Venant model. Two of them also have adverse distributions that bring worse flooding.
@@
 It is usually seen that uncertainty in $D$ makes very little difference.
-See for example Iooss and Lemaître (2015).
+See for example Iooss & Lemaître (2015).
 Therefore we fix $D=5000$.
@@
 With probability $0.01$ both $Q$ and $K_{s}$ are adverse.
-The distribution of $K_{s}$ is customarily given as $\mathcal{N}(30,7.5^{2})$ and then truncated to a range such as $(0,\infty)$ in Limbourg and De Rocquigny (2010) or $(15,\infty)$ in Iooss (2011).
+The distribution of $K_{s}$ is customarily given as $\mathcal{N}(30,7.5^{2})$ and then truncated to a range such as $(0,\infty)$ in Limbourg & De Rocquigny (2010) or $(15,\infty)$ in Iooss (2011).
 We have chosen a nominal Gamma distribution, so we do not need any truncation.
@@
 The distribution of $Q$ is customarily given as Gumbel with mode $1013$ and scale $558$.
-This was found by a maximum likelihood fit, reported in Limbourg and De Rocquigny (2010).
+This was found by a maximum likelihood fit, reported in Limbourg & De Rocquigny (2010).
 The Gumbel distribution has an extreme value justification, but it also gives negative values which are unphysical.
@@
 Figure 3: We plot variance versus sample size for $\hat{\mu}_{\mathrm{MC}}$, $\hat{\mu}_{\mathrm{RQMC}}$, $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$, $\hat{\mu}_{\mathrm{RQMC}(2)}$ and $\hat{\mu}_{\mathrm{RQMC}(L)}$ for the Saint-Venant flood depth (15).
+The top to bottom ordering of the curves in the legend is the same as they have at $n=4096$.
 The sample size allocations for $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$, $\hat{\mu}_{\mathrm{RQMC}(2)}$, and $\hat{\mu}_{\mathrm{RQMC}(L)}$ used $\rho=2$.
@@
 In Figure 4 we consider integer allocations with different values of $\rho$ using equation (9) to select $n_{\ell}$.
-For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope close to $-1.8$.
+For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope close to 1.8-1.8.
 The curve for $\rho=2$ which is also in Figure 3 is best at larger $n$.
-The case of $\rho=1$ has proportional allocation and may have an empty stratum for $n\leqslant 128$ (as $128*0.01<2$).
-Figure 4: Variance of $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$ versus the sample size $n$ for $\rho\in\{1,2,3,+\infty\}$. The reference line has slope $-1.8$.
+The case of $\rho=1$ has proportional allocation and may have an empty stratum for $n\leqslant 128$ (as 1280.01<2128*0.01<2).
+Figure 4: Variance of $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$ versus the sample size $n$ for $\rho\in\{1,2,3,+\infty\}$. The reference line has slope 1.8-1.8.
 8 Discussion
@@
 We speculate that the negative correlation advantage, if it is present, is greatest when several $\beta_{\ell}$ are large.
-We also tried a complex importance sampling problem from Hesterberg (1995) modeling reliability of oil inventory for an energy utility.
-The integrand was of the mixture importance sampling form (1).
+For some problems, the integrand is a simple function within each stratum.
+A more interesting expression arises in mixture importance sampling.
+There we estimate $\mu=\int g(\bm{x})p(\bm{x})\mathrm{\,d}\bm{x}$ for a probability density function $p$ by
+$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\frac{g(\bm{x}_{i})p(\bm{x}_{i})}{\sum_{\ell=1}^{L}\alpha_{\ell}p_{\ell}(\bm{x}_{i})}$ (16)
+for $\bm{x}_{i}\sim\sum_{\ell=1}^{L}\alpha_{\ell}p_{\ell}$.
+We tried a complex importance sampling problem from Hesterberg (1995) modeling reliability of oil inventory for an energy utility.
+The estimator was of the mixture importance sampling form (16).
+An interesting aspect of mixture importance sampling is that it benefits from using $p_{\ell^{\prime}}(\bm{x}_{i})$ when $\bm{x}_{i}$ was drawn from $p_{\ell}$ for $\ell\neq\ell^{\prime}$.
 We found only a moderate improvement with RQMC giving roughly $O(n^{-1.15})$ variance.
-The example comes from Hesterberg (1995) and was also studied by Li et al.
+The example was also studied by Li et al.
 (2013).
-Here we only mention it to show that RQMC with mixture sampling will not always give an improvement that is practically meaningful.
-The relevant integrands were discontinuous and we speculate that they may have had important high order interactions.
+Here we only mention it to show that RQMC with mixture sampling will not always give an an extremely large improvement.
+The relevant integrands were discontinuous and we speculate that they may also have had important high order interactions.
 We might also have a variable with a mixed distribution including a continuous component with a density function and $L$ atoms.
@@
 We have considered one categorical variable.
-Some problems may well have two or more categorical variables.
+Some problems will have two or more categorical variables.
 Suppose that there are $K$ categorical variables where the $k$’th one has $L_{k}$ levels.
@@
 Thanks to Tim Hesterberg and Wentao Li for help with the oil inventory problem.
-This work was supported by the National Science Foundation under grant DMS-2152780.
-Appendix A Monotonocity of $I_{0}$
-Here we prove Proposition 4.
+We thank two anonymous referees and an associate editor for comments that helped us improve the paper.
+[…]
+This appendix contains the proofs of our propositions.
+[…]
+Proof.
+Write $A=r/n+a$ and $B=s/n+b$ for integers $0\leqslant r\leqslant s<n$ and $a,b\in[0,1/n)$.
+Then $\beta=t/n+c$ for t=srt=s-r and 1/n<c<1/n-1/n<c<1/n.
+If $r=s$, then $0\leqslant n_{*}\leqslant 1$ and then (5) holds trivially.
+For $s>r$ we may write
+$A,B)=[A,(r+1)/n)\,\cup\,[(r+1)/n,s/n)\,\cup\,[s/n,s/n+b).$
+Then $[A,B)$ contains at least sr1s-r-1 of the intervals $I_{k}=[(k-1)/n,k/n)$ for $k=1,\dots,n$ and it intersects with at most sr+1s-r+1 of them.
+Therefore $|n_{*}-(s-r)|\leqslant 1$.
+Now $\lceil n\beta\rceil=s-r+\lceil nc\rceil$ and so $s-r\geqslant\lceil n\beta\rceil-1$ establishing the lower limit in (5).
+Similarly $\lfloor n\beta\rfloor=s-r+\lfloor nc\rfloor$ so $s-r\leqslant\lfloor n\beta\rfloor+1$ establishing the upper limit in (5).
+∎
+[…]
+Proof.
+For $S_{\ell}$ defined at (3), sample $i\in S_{\ell}$ holds if and only if $B_{\ell-1}\leqslant z_{i1}<B_{\ell}$ or equivalently $\sum_{1\leqslant r<\ell}2^{-\kappa_{r}}\leqslant z_{i1}<\sum_{1\leqslant r\leqslant\ell}2^{-\kappa_{r}}$ which we can write as
+$\frac{\gamma}{2^{\kappa_{\ell}}}\leqslant z_{i1}<\frac{\gamma+1}{2^{\kappa_{\ell}}}$ (17)
+for integer $\gamma=\sum_{1\leqslant r<\ell}2^{\kappa_{\ell}-\kappa_{r}}\geqslant 0$.
+Because the net is stratified, the number of observations $i$ that satisfy equation (17) is $2^{m-\kappa_{\ell}}=2^{m_{\ell}}$.
+Then $\{u_{i}\mid i\in S_{\ell}\}\subset[0,1)^{s}$ has cardinality $2^{m_{\ell}}$ making it a $(t_{\ell},m_{\ell},s)$-net in base $2$ for some value $t_{\ell}\leqslant m_{\ell}$.
+Now consider a dyadic box in $0,1]^{s}$ of the form
+$\prod_{j=1}^{s}\Bigl[\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr).$
+This box has volume $2^{-|\bm{k}|}$. The number of points $\bm{u}_{i}$ from stratum $\ell$ in this box equals the number of points $\bm{z}_{i}\in[0,1)^{s+1}$ that belong to
+$\Bigl[\frac{\gamma}{2^{\kappa_{\ell}}},\frac{\gamma+1}{2^{\kappa_{\ell}}}\Bigr)\times\prod_{j=1}^{s}\Bigl[\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr).$
+If $m-\kappa_{\ell}-|\bm{k}|\geqslant t$, then by the digital net property of $\bm{z}_{1},\dots,\bm{z}_{n}$, this number is $2^{m-\kappa_{\ell}-|\bm{k}|}$.
+That is, if $m_{\ell}\geqslant|\bm{k}|+t$, then the number of such points is $2^{m_{\ell}-|\bm{k}|}$.
+∎
+[…]
+Proof.
+The local discrepancy of $\bm{z}_{1},\dots,\bm{z}_{n}$ at the point $(b,\bm{c})\in[0,1]^{s+1}$ is
+$\delta(b,\bm{c})=\frac{1}{n}\sum_{i=1}^{n}\bm{1}\{0\leqslant v_{i}<b\}\bm{1}\{\bm{z}_{i}\in[\bm{0},\bm{c})\}-b\times\mathrm{Vol}([\bm{0},\bm{c}))$
+where $\mathrm{Vol}([\bm{0},\bm{c}))=\prod_{j=1}^{s}c_{j}$. Because $\bm{z}_{1},\dots,\bm{z}_{n}$ are a digital net, $|\delta(b,\bm{c})|=O(n^{-1}\log(n)^{s})$.
+The sample points for stratum $\ell$ are the $n_{\ell}$ points $\bm{x}_{i}$ for which $B_{\ell-1}\leqslant v_{i}<B_{\ell}$.
+By equation (5), $n_{\ell}=n\beta_{\ell}+O(1)$.
+The local discrepancy within stratum $\ell$ at $\bm{c}\in[0,1)^{s}$ is
+$\ \quad\frac{1}{n_{\ell}}\sum_{i=1}^{n}\bm{1}\{B_{\ell-1}\leqslant v_{i}<B_{\ell}\}\bm{1}\{\bm{x}_{i}\in[\bm{0},\bm{c})\}-\mathrm{Vol}([\bm{0},\bm{c}))$
+$=\frac{n}{n_{\ell}}\Bigl(\beta_{\ell}\mathrm{Vol}([\bm{0},\bm{c}))+\delta(B_{\ell},\bm{c})-\delta(B_{\ell-1},\bm{c})\Bigr)-\mathrm{Vol}([\bm{0},\bm{c})).$
+Its absolute value is at most
+$\Bigl|\frac{n\beta_{\ell}}{n_{\ell}}-1\Bigr|\mathrm{Vol}([\bm{0},\bm{c}))+\frac{2n}{n_{\ell}}O(n^{-1}\log(n)^{s})$ $=O(n_{\ell}^{-1}\log(n_{\ell}/\beta_{\ell})^{s})$
+$=O(n_{\ell}^{-1}\log(n_{\ell})^{s})$
+and so the star discrepancy of $\bm{u}_{i}$ for $i\in S_{\ell}$ is $O(n_{\ell}^{-1}\log(n_{\ell})^{s})$. ∎
+[…]
 Proof.
@@
 Then we can write $\tilde{I}_{0}(\gamma\!\mid\!\rho)=h(x)$ where
-$h(x)$ $=\biggl{(}\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2-2\rho/(\rho+1)-\rho x}\biggr{)}\biggl{(}\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr{)}^{\rho}$
-$=\biggl{(}\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2/(\rho+1)-\rho x}\biggr{)}\biggl{(}\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr{)}^{\rho}.$
+$h(x)$ $=\biggl(\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2-2\rho/(\rho+1)-\rho x}\biggr)\biggl(\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr)^{\rho}=\biggl(\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2/(\rho+1)-\rho x}\biggr)\biggl(\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr)^{\rho}.$
 It is enough to show that $h(x)$ is non-decreasing as $x$ either increases from zero or decreases from zero. We will do this by showing that $xh^{\prime}(x)\geqslant 0$.
@@
 $S_{-}(r)$ $=S_{-}(r,x)=\sum_{\ell=1}^{L}\alpha_{\ell}^{2/(\rho+1)-\rho x}\log(\alpha_{\ell})^{r}.$
-Their derivatives with respect to $x$ are $S_{+}(r+1)$ and $-\rho S_{-}(r+1)$, respectively. Also, $S_{+}(r,0)=S_{-}(r,0)$ which we write as $S(r)$.
-Now we write
-$h(x)=S_{-}(0)S_{+}(0)^{\rho}$
-so
-$h^{\prime}(x)=-\rho S_{-}(1)S_{+}(0)^{\rho}+\rho S_{-}(0)S_{+}(0)^{\rho-1}S_{+}(1).$
-Then
+Their derivatives with respect to $x$ are $S_{+}(r+1)$ and $\rho S_{-}(r+1)$, respectively. Also, $S_{+}(r,0)=S_{-}(r,0)$ which we write as $S(r)$.
+Now we write $h(x)=S_{-}(0)S_{+}(0)^{\rho}$ so $h^{\prime}(x)=-\rho S_{-}(1)S_{+}(0)^{\rho}+\rho S_{-}(0)S_{+}(0)^{\rho-1}S_{+}(1)$. Then
 $\frac{h^{\prime}(x)}{\rho S_{+}(0)^{\rho-1}}$ $=S_{+}(1)S_{-}(0)-S_{+}(0)S_{-}(1)$
@@
 as required. ∎
+[…]
+Proof.
+From equation (3.2)
+$I_{0}(1\!\mid\!2)=\frac{L}{\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}^{2/3}\bigr)^{3}}\quad\text{and}\quad I_{0}(3\!\mid\!2)=\frac{\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}^{1/2}\bigr)^{2}}{\bigl(\sum_{\ell=1}^{L}\alpha_{\ell}^{2/3}\bigr)^{3}}.$
+These expressions have the same denominator and then the result follows from the Cauchy-Schwarz inequality applied to the numerators. ∎
+[…]
+Proof.
+First, if $n^{\prime}>0$ then we can show that there is an ‘eligible’ index $\ell$ with $n_{\ell}\leqslant n^{\prime}$, that can be doubled without causing $\sum_{\ell=1}^{L}n_{\ell}>n$.
+At any stage in the algorithm, $n_{\ell^{*}}$ for $\ell^{*}=\operatorname*{argmin}_{1\leqslant\ell\leqslant L}n_{\ell}$ is a power of two.
+It divides $n$ and it divides every $n_{\ell}$, so it also divides the nonzero value $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$, hence $n_{\ell^{*}}\leqslant n^{\prime}$.
+Finally, every step with $n^{\prime}>0$ reduces $n^{\prime}$, and so $n^{\prime}$ must reach zero in a finite number of steps starting from $n^{\prime}=n-L$.
+∎
+[…]
+Proof.
+The same argument applies to both criteria.
+We prove it for criterion 1.
+Fixing $\bm{n}$ and $\tilde{\bm{n}}$ we see that the maximum over $\bm{\tau}$ of $R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau})$ arises when $\tau_{\ell}=0$ for all $\ell$ except $\ell^{*}(\bm{n},\tilde{\bm{n}})=\operatorname*{argmax}_{\ell}(\tilde{n}_{\ell}/n_{\ell})^{\rho/2}$ with $\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}$ then taking any value $\lambda>0$.
+Now, we can go back to the initial problem and rewrite (13) as:
+$\min_{\bm{n}\in\Delta_{N}}\max_{\tilde{\bm{n}}\in\Delta_{N}}\Bigl(\frac{\tilde{n}_{\ell^{*}(\bm{n},\tilde{\bm{n}})}}{{n}_{\ell^{*}(\bm{n},\tilde{\bm{n}})}}\Bigr)^{\rho/2}$
+after noting that the factor $\alpha_{\ell^{*}(\bm{n},\tilde{\bm{n}})}\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}^{1/2}$ cancels from numerator and denominator.
+To maximize this ratio over $\tilde{\bm{n}}$ one takes $\tilde{\bm{n}}$ to be the vector with value NL+1N-L+1 in component $\ell^{*}=\operatorname*{argmin}_{\ell}n_{\ell}$ and ones elsewhere.
+Then the minimax problem becomes $\min_{\bm{n}\in\Delta_{N}}\bigl((N-L+1)/\min_{1\leqslant\ell\leqslant N}n_{\ell}\bigr)^{\rho/2}$.
+The solution is to maximize the minimum value of $n_{\ell}$.
+We do that by making $\min_{\ell}n_{\ell}=\lfloor N/L\rfloor$.
+∎
+[…]
+Our model for sample allocation is that we start with $n_{\ell}=1$ for $\ell=1,\dots,L$.
+Then we repeatedly increment $n_{\ell^{*}}$ by one for $\ell^{*}=\arg\max_{\ell}\alpha_{\ell}^{2}(n_{\ell}^{-\rho_{\ell}}-(n_{\ell}+1)^{-\rho_{\ell}})$ because this gives the stratum that most reduces the criterion.
+For our RQMC estimate to be consistent, we require $\min_{\ell}n_{\ell}\to\infty$, and so we assume that all of the $n_{\ell}$ are large enough to allow a Taylor approximation below.
+We also ignore ties in this analysis.
+Using $(n_{\ell}+1)^{-\rho_{\ell}}=n_{\ell}^{-\rho_{\ell}}-\rho_{\ell}n_{\ell}^{-(\rho_{\ell}+1)}+O(n_{\ell}^{-\rho_{\ell}-2})$, we simplify the selection rule to $\ell^{*}=\arg\max_{\ell}\alpha_{\ell}^{2}\rho_{\ell}/n_{\ell}^{\rho_{\ell}+1}.$ Then for $1\leqslant\rho_{\ell}<\rho_{\ell^{\prime}}\leqslant 3$ under our greedy allocation $n_{\ell^{\prime}}$ will only get a new point if
+$n_{\ell^{\prime}}<\Bigl(\frac{\alpha_{\ell^{\prime}}^{2}\rho_{\ell^{\prime}}}{\alpha_{\ell}^{2}\rho_{\ell}}n_{\ell}^{\rho_{\ell}+1}\Bigr)^{1/(\rho_{\ell^{\prime}}+1)}$
+and so we expect $n_{\ell^{\prime}}\lesssim n_{\ell}^{(\rho_{\ell}+1)/(\rho_{\ell^{\prime}}+1)}$, where $\lesssim$ denotes inequality up to a multiplicative constant.
+For the most extreme case, with $\rho_{\ell}=1$ and $\rho_{\ell^{\prime}}=3$, we get $n_{\ell^{\prime}}\lesssim n_{\ell}^{1/2}$.